### v1

# Chen, Simchi-Levi, and Wang

[…]
Chen, Simchi-Levi, and Wang
[…]
Experimenting on MDPs
[…]
Experimenting on Markov Decision Processes with Local Treatments
[…]
As service systems grow increasingly complex and dynamic, many interventions become localized, available and taking effect only in specific states. This paper investigates experiments with local treatments on a widely-used class of dynamic models, Markov Decision Processes (MDPs). Particularly, we focus on utilizing the local structure to improve the inference efficiency of the average treatment effect. We begin by demonstrating the efficiency of classical inference methods, including model-based estimation and temporal difference learning under a fixed policy, as well as classical A/B testing with general treatments. We then introduce a variance reduction technique that exploits the local treatment structure by sharing information for states unaffected by the treatment policy. Our new estimator effectively overcomes the variance lower bound for general treatments while matching the more stringent lower bound incorporating the local treatment structure. Furthermore, our estimator can optimally achieve a linear reduction with the number of test arms for a major part of the variance. Finally, we explore scenarios with perfect knowledge of the control arm and design estimators that further improve inference efficiency.
[…]
Experimentation, A/B Testing, Markov Decision Process, Local Treatment
1 Introduction
Randomized experiments, commonly known as A/B testing, have emerged as an indispensable tool, enabling organizations across various sectors—including service, retail, manufacturing, and healthcare—to empirically test and refine their strategies for enhancing customer engagement and satisfaction (Imbens and Rubin 2015, Kohavi and Thomke 2017, Kohavi et al. 2020, Koning et al. 2022).
However, as service systems become more complex and dynamic, the implementation and analysis of randomized experiments present more challenges, as well as new opportunities. One notable development in these dynamical systems is the increasing prevalence of localized treatments or policies that target specific customer interactions based on recent behaviors or other contextual factors (i.e., state). For instance, an online marketplace or service platform might offer a “welcome-back” coupon to the customer only when she has not engaged with the service for an extended period (e.g., Sephora’s Welcome Back Event Dooley 2020). Similarly, physical retail locations like pharmacies and markets may provide incentives like cash back to the customer only when she makes in-store purchases (e.g., Bed Bath & Beyond Mohammed 2015). Likewise, a typical intervention for audio stream recommender systems may involve modifying banner promotions alone, without altering other system components throughout the iterative recommendation process (e.g., Spotify Maystre et al. 2023). These localized treatments are strategic, usually aiming to optimize the lifetime value of customers by tailoring experiences and offers to individual needs and histories.
To analyze the effectiveness of local treatments in dynamic systems, we employ Markov Decision Processes (MDPs), which are commonly used for modeling complex scenarios including customer relationships (Pfeifer and Carraway 2000), inventory control (Gong and Simchi-Levi 2023), two-sided platforms (Johari et al. 2022), healthcare (Yu et al. 2021) and many others. Despite the broad applications of MDPs, experimentation on MDPs remains relatively unexplored. Existing research typically models global treatments applied uniformly across all states of an MDP. However, this approach, which ignores the potential structure of the treatment, may lead to prohibitively large variances. For example, Farias et al. (2022) demonstrate that the lower bound for any unbiased estimators can sometimes grow exponentially with the number of states. They further propose reducing the variance by sacrificing unbiasness. Other prior work, such as Shi et al. (2020), Syrgkanis and Zhan (2023) rely heavily on function approximations that may introduce additional bias, especially when the model is mis-specified. Alternatively, based on the practical observations of local treatments, in this paper we elaborate on specific treatment structures and raise the following research question:
Whether and how the local structure of treatments can be utilized to improve the inference efficiency on MDPs?
In this paper, we systematically explore the potential of leveraging the structure of local treatments to optimize data utilization from A/B testing within an MDP environment. Our findings conclusively answer this question in an affirmative and optimal manner.
1.1 An illustrative Example
To be more concrete about our research question, we further introduce an application example on improving customer lifetime value, adapted from a seminal work in the marketing literature by Pfeifer and Carraway (2000) and closely related to customer retention and churn management (Lemmens and Gupta 2020, Kanoria et al. 2024). This example will be revisited throughout the paper.
Example 1.1 (Customer Lifetime Value)
Drawing from the classic Recency, Frequency, Monetary value (RFM) framework, it is often posited by companies that the probability a customer will purchase at the end of any period is solely a function of her recency—the number of periods since her last purchase. Pfeifer and Carraway (2000) model recency as the state space of a MDP and characterizes customer lifetime value using value functions. The state transitions occur as follows: each purchase resets the customer’s recency to one; conversely, if no purchase is made, the customer’s recency increases by one period. When a customer’s recency is high (for example, $\geq$5), Pfeifer and Carraway (2000) aggregate these instances into a single state, denoted as $s^{5}$, and regard it as representing an inactive customer.
Currently, in order to encourage repeat business and build customer loyalty, the company is exploring the impact of issuing coupons to customers with high recency (i.e., inactive customers at state $s^{5}$) to attract them back. The coupon treatment is considered local, affecting only the state transitions and immediate reward at state $s^{5}$, while leaving all other states unchanged before and after the treatment. Our study can then enable the company to assess how such a coupon policy might affect customer lifetime value in the long run.
1.2 Overview of Main Results
To demonstrate our results, we start with what is arguably the most fundamental and illustrative scenario: the new local treatment available in only one specific state. We refer to such a local treatment as Single-State Treatment (SST). In this setup, the treatment policy of interest applies the new SST whenever possible and adheres to the control policy in all other states. Our objective is to estimate and infer the Average Treatment Effect (ATE), which quantifies the difference in long-term cumulative rewards between the treatment policy and the control policy. We summarize and highlight our main results in detail below.
Efficient Inference under a Fixed Policy. Instead of simultaneously evaluating both the treatment and control policies, we initially focus on presenting estimation methods for the long-term cumulative rewards of a single, specific policy, say, either the treatment or the control.
The two principal approaches used in the current literature are model-based estimation and temporal difference (TD) learning. Model-based estimation involves directly estimating all parameters of the MDP under a fixed policy to subsequently compute long-term rewards. Conversely, TD learning is more straightforwards to fit the long-term rewards according to the Bellman equation. Despite their distinct approaches, we establish that both methods are equivalent and demonstrate asymptotic normality. This means that the two estimators yield identical results and share confidence intervals. This equivalence allows the theoretical results derived from one method to be applicable to the other, providing the flexibility to select the method that is simpler to analyze. Moreover, we derive a variance lower bound for all unbiased estimators using a constrained Cramér-Rao bound. This highlights the inherent challenges of the inference task and confirms that both methods are asymptotically efficient compared to all other unbiased estimators, which may be of independent interest for the reinforcement learning (RL) literature.
Variance Reduction Utilizing the Local Treatment Structure. To accurately infer the ATE, we simultaneously estimate the long-term cumulative rewards for both the treatment and control policies. Particularly, we follow the spirit of A/B testing and apply the treatment and control policies randomly at each time period with equal probability. Significant challenges arise from the fact that the data collected are generated by a mixed policy—neither solely the treatment nor the control policy—and our focus is on the long-term cumulative rewards, not just instant rewards. We first establish an unbiased estimator for the ATE, treating the treatment and control policies as completely distinct, without leveraging the SST structure. This estimator is shown to be asymptotically normal, and its variance is almost optimal when the SST structure is not used, as confirmed by our derived variance lower bound.
More importantly, we introduce an estimator that incorporates the SST structure by sharing data collected from states where the new treatment is not applied. This estimator can always achieve a smaller variance, demonstrating the benefits of utilizing the SST structure. Importantly, when expanding to multiple test arms, we illustrate that a major part of the asymptotic variance can be reduced linearly with the number of test arms compared to the estimator that does not use the SST structure. Additionally, we derive another variance lower bound for all unbiased estimators even with the knowledge of the SST, validating the optimality of our proposed estimator.
Further Variance Reduction with Perfect Information about the Control Arm. In many applications, the control policy has been operational as the default setting within systems for an extended period, yielding a wealth of offline data. Inspired by this abundance of historical information, we explore a scenario where we possess perfect knowledge of the control policy, including its state transition probabilities and immediate rewards. In response, we introduce a new estimator that is asymptotically normal, leveraging both our comprehensive understanding of the control policy and the SST structure. Our findings demonstrate that this approach can further reduce the asymptotic variance, enhancing the estimator’s effectiveness.
Having demonstrated the variance reduction achieved through SST, we also expand our framework and generalize the results to encompass more general local treatments. This extension implies that the new treatment may be applicable to multiple states, broadening the scope and applicability of our results.
Moreover, we reveal that under SST, the treatment effect remains consistently positive (or negative) for all states before and after the intervention. This indicates that if the new treatment benefits one state, it benefits all others as well. Also, we discuss that our information sharing concept is orthogonal to and compatible with the bias-variance trade-offs (e.g., Farias et al. 2022) and function approximation approaches (e.g., Shi et al. 2020, Syrgkanis and Zhan 2023) developed for the global treatments. This means that many developments under the global treatments can be further improved with the presence of the local treatments by our information sharing idea.
1.3 Related Works
Structure-aware Experiment Design. Experimental design originated in economics with basically model-free methods. In recent years, researchers in operations research and management science instead realize that there are many useful structures in operations scenarios to understand, analyze and improve experimentation. Bojinov et al. (2023), Hu and Wager (2022), Ni et al. (2023), Jia et al. (2023), Xiong et al. (2024) consider the switchback design under various scenarios, from temporal interference to spatial interference, from stationary environment to non-stationary environment, and from theory driven perspective to data driven perspective. Glynn et al. (2020), Munro et al. (2024), Wager and Kuang (2021), Bajari et al. (2021), Johari et al. (2022), Zhu et al. (2024) study the experiments happening on two-sided markets and a common practice is to model the system as a continuous-time Markov chain. Li et al. (2015) leverage the sparsity structure of cross-product demand elasticities to greatly reduce the number of experiments required. Li et al. (2023) carefully take the queuing structure into account and propose efficient estimators under the stochastic congestion. Xiong et al. (2023) investigate the setting with multiple units where the starting time of the treatment may vary by unit. Zhan et al. (2021, 2023) study the post-experiment inference for contextual bandits, which does not capture the state transition as our MDPs. Simester et al. (2020) pioneer at considering treatment structures and suggest comparing outcomes only for customers exposed to different treatment and control actions to improve efficiency empirically. Compared to theirs, we further incorporate complex transition dynamics and provide theoretical guarantees of estimations. The long-term effect with the long-term treatment under the short-term measurement is recently studied by Huang et al. (2023). This paradigm of structure-aware experiment design has witnessed great success in theory and practice in above works with either significant reduction of bias or substantial improvement of efficiency.
Experimenting on MDPs with General Treatments. Experimenting on MDPs is more relevant to our work. Many works adopt MDPs to capture the dynamical system and investigate the long-term outcomes (Pfeifer and Carraway 2000, Murphy 2003, Simester et al. 2006, Tran et al. 2023, Shi et al. 2020, 2023, Farias et al. 2022, 2023), where most previous works consider general treatment, i.e. a treatment will be posed to every state of MDP. The seminal work by Farias et al. (2022) studies the average reward setting. They prove the first asymptotic lower bound for unbiased inference on MDPs and show that this lower bound can sometimes grow exponentially with the number of states. They also construct an LSTD(0)-type OPE estimator that is efficient and achieves this lower bound. In our work, we show that similar phenomena can also be observed in the discounted reward settings. Moreover, Farias et al. (2022) propose an alternative estimator that achieves a notable bias-variance trade-off with an upper bound of asymptotic variance. In contrast, all estimators discussed in this paper are asymptotically unbiased with exact formulae of asymptotic variance preferable for accurately determining the local treatment effect. We instead manage to reduce variance by taking into account the treatment structure. We also discuss the adaptation of their bias-variance trade-off estimator to the discounted reward setting (Farias et al. 2023) in Section 9.2. Furthermore, given the variance lower bound, some recent work like Shi et al. (2023), Zhang et al. (2023) considers asymptotic results with linear function approximation. Note that our methods are orthogonal to the bias-variance trade-off estimator and the function approximation method, and thus are compatible with them.
In addition, there is another line of work including Syrgkanis and Zhan (2023), Cheikhi and Russo (2023), Shi et al. (2020), Khamaru et al. (2021), Tran et al. (2023) that considers asymptotic inference on MDPs but under episodic setting or with the assumption of a generative model, which means the MDP can be independently simulated many times. In contrast, we consider estimations that can only sample one trajectory from a non-episodic MDP, which restricts our access to the data and thus imposes additional challenges.
1.4 Structures and Notations
In Section 2, we formally introduce the model and the formulation, introducing the MDP, the treatment policy and the control policy, and the average treatment effect that we care about. Section 3 presents the estimation methods for the long-term cumulative rewards of a single, specific policy. In Section 4, we discuss the idea of classical A/B testing in our setting and how to estimate the long-term cumulative rewards for the treatment and control policies from A/B testing. Section 5 provides a new estimator based on the idea of information sharing for the local treatment. Section 6 studies the setting with perfect information of control. Section 7 generalizes the analysis to the general local treatments. In Section 8, we provide some numerical results verifying the idea of information sharing. Section 9 covers the discussion on some useful properties of the local treatment, and the compatibility with other popular techniques in RL. We conclude our paper in Section 10.
We denote for any $N\in\mathbb{Z}^{+},[N]=\{1,2,\cdots,N\}$. For any vector $x$ and matrix ${\bm{A}}$, we define the semi-norm induced by the matrix $A$ as $\|x\|_{\bm{A}}\coloneqq\sqrt{x^{\top}{\bm{A}}x}$. Additionally, for any function on the state space $f\colon{\mathcal{S}}\to\mathbb{R}$, we denote $\operatorname{Diag}_{s\in{\mathcal{S}}}(f(s))\coloneqq\operatorname{Diag}(f(s^{1}),\cdots,f(s^{K}))\in\mathbb{R}^{K\times K}$. Moreover, we define $a\vee b\coloneqq\max\{a,b\}$ and $a\land b\coloneqq\min\{a,b\}$.
2 Model
We start with the general definition of experimenting on MDPs and then specify the Single-State Treatment (SST), which serves as a simple yet illustrative example of local treatment. In Section 7, we generalize this concept to broader local treatment scenarios.
Consider a (non-episodic, finite-state, infinite-horizon) MDP with state space ${\mathcal{S}}:=\{s^{i}\}_{i\in[K]}$ with cardinality $K$. At time $t\in\mathbb{N}$, we denote the state of the MDP by $s_{t}\in{\mathcal{S}}$. Associated with each state is a set of available actions, ${\mathcal{A}}$, which influence the state transitions through the transition probability function $P:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\rightarrow[0,1]$. Here, the next state $s_{t+1}$ follows the distribution $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$.
For each state-action pair $(s,a)$, the instant reward $r$ earned is independently generated by a Gaussian distribution [^1] With standard arguments, our results can be extended to the setting where the second moment of the noise is finite., i.e., $r\sim{\mathcal{N}}(r(s,a),\sigma_{r}^{2}(s,a))$, where we assume $\sigma^{2}_{r}(s,a)<\infty$ for all $(s,a)$ pairs.
A policy $\pi:{\mathcal{S}}\rightarrow{\mathcal{A}}$ maps states to (random) actions. The value function is defined as the cumulative discounted reward starting from state $s\in{\mathcal{S}}$ under a policy $\pi$,
$V^{\pi}(s)\coloneqq\mathbb{E}\left[{\sum_{i=0}^{\infty}\gamma^{i}r(s_{i},a_{i})\mid\pi,s_{0}=s}\right],$ (1)
where $10\leq\gamma<$ 1 is a discount factor that reflects the trade-off between instant and future rewards. The value function has been widely applied to quantify long-term outcomes in various fields including marketing (Pfeifer and Carraway 2000, Simester et al. 2006), ride-hailing platforms (Shi et al. 2023), and video streaming service (Tran et al. 2023, Maystre et al. 2023). We further use ${\bm{V}}^{\pi}\in\mathbb{R}^{K}\coloneqq(V^{\pi}(s^{1}),\cdots,V^{\pi}(s^{K}))^{\top}$ to denote value functions of all states. Moreover, associated with each $\pi$, ${\bm{r}}^{\pi}\in\mathbb{R}^{K}$ is the expected reward under each state under policy $\pi$, defined as ${\bm{r}}^{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}r(s,a)\pi(a\mid s)$ for $s\in{\mathcal{S}}$. ${\bm{P}}^{\pi}\in\mathbb{R}^{K\times K}$ is the transition matrix under $\pi$ with ${\bm{P}}^{\pi}(s,s^{\prime})\coloneqq\sum_{a\in{\mathcal{A}}}P(s^{\prime}\mid s,a)\pi(a\mid s)$ for $s,s^{\prime}\in{\mathcal{S}}$.
We then introduce a special class of treatment policies that we call the Single-State Treatment (SST). Specifically, there exists a pre-determined “crucial” state, which without loss of generality, we assume to be $s^{1}$. The treatment of interest is applied exclusively when the system’s state, $s_{t}$, coincides with $s^{1}$.
Formally, we assume that ${\mathcal{A}}=\{{\mathfrak{c}},{\mathfrak{t}}\}$ regardless of states, while the action ${\mathfrak{t}}$, representing the new treatment, is only applicable in the crucial state $s^{1}$, while ${\mathfrak{c}}$ denotes the control or default action. This restriction ensures that the treatment is targeted and its effects are isolated to the crucial state.
Building on the definition of SST, we next define two policies and the corresponding treatment effect we aim to assess.
  • The Control Policy $\pi^{\mathfrak{c}}$: This policy does not implement the new intervention, maintaining $\pi^{\mathfrak{c}}(s)\equiv{\mathfrak{c}}$ across all states $s\in{\mathcal{S}}$. This is typically the default or existing policy used by the organization. Denote the associated transition matrix as ${\bm{P}}^{\mathfrak{c}}$, where the entries are exactly $P(\cdot\mid\cdot,{\mathfrak{c}})$. The vector ${\bm{r}}^{\mathfrak{c}}\in\mathbb{R}^{K}$ is the unknown mean reward for each state under the control whose entries are just $r(\cdot,{\mathfrak{c}})$.
  • The Treatment Policy $\pi^{\mathfrak{t}}$: This policy implements the new intervention ${\mathfrak{t}}$ exclusively whenever the state is $s^{1}$, while defaulting to ${\mathfrak{c}}$ in all the other states. Formally, $\pi^{\mathfrak{t}}(s^{1})\equiv{\mathfrak{t}}$ and $\pi^{\mathfrak{t}}(s)\equiv{\mathfrak{c}}$ for all $s\in{\mathcal{S}}\setminus s^{1}$. Accordingly, we denote the associated transition matrix and the mean reward vector as ${\bm{P}}^{\mathfrak{t}}$ and ${\bm{r}}^{\mathfrak{t}}$, respectively.
The treatment effect that we are interested in is the difference of the value function under the treatment policy and the control policy. For simplicity, we specially abbreviate the value functions under $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$, $V^{\pi^{\mathfrak{t}}}(s)$ and $V^{\pi^{\mathfrak{c}}}(s)$, as $V^{{\mathfrak{t}}}(s)$ and $V^{{\mathfrak{c}}}(s)$, respectively. Specifically, for a given initial state $s$, the estimand of our interest is the average treatment effect (ATE),
$\Delta(s)=V^{{\mathfrak{t}}}(s)-V^{{\mathfrak{c}}}(s).$
For all the states, we write the treatment effect in a more compact way as ${\bm{\Delta}}\coloneqq{\bm{V}}^{{\mathfrak{t}}}-{\bm{V}}^{{\mathfrak{c}}}$, where ${\bm{V}}^{{\mathfrak{t}}}$ and ${\bm{V}}^{{\mathfrak{c}}}$ are abbreviations of ${\bm{V}}^{\pi^{\mathfrak{t}}}$ and ${\bm{V}}^{\pi^{\mathfrak{c}}}$, respectively. The central research question of our paper is to estimate the ATE, and especially, how the SST structure can be taken into account and utilized.
Before we conclude the session, we would like to introduce an important measure, which plays an important role in our theoretical results and is also commonly used in the RL literature (Agarwal et al. 2022).
Definition 2.1 (Discounted State-Visitation Measure)
Given a policy $\pi$, for any two states $s,s^{\prime}$, the Discounted State-Visitation Measure is the infinite discounted summation of the probability of visiting state $s$ starting from state $s^{\prime}$, i.e.,
$\rho^{\pi}_{s^{\prime}}(s)$ $\coloneqq\sum_{i=0}^{\infty}\gamma^{i}\mathbb{P}^{\pi}(s_{i}=s\mid s_{0}=s^{\prime})\,.$
Intuitively, the discounted state-visitation measure quantifies the likelihood that policy $\pi$ will lead to state $s$ in the future, starting from the initial state $s^{\prime}$. When the discounted factor $\gamma$ approaches $1$, this measure closely approximates the expected number of times the state $s$ is visited. This metric effectively captures the underlying dynamics of the MDP. Given that our intervention is exclusively applied to state $s^{1}$, the frequency with which state $s^{1}$ is visited becomes a crucial factor in assessing the impact of the treatment.
3 Preliminaries: Inference under A Fixed Policy
Before presenting the methods for experimenting with and comparing different policies, in this section, we first discuss how to perform inference on the value function under a fixed policy $\pi$, as an important cornerstone. Estimating the value function within MDPs is not a new task in the literature. There are two streams of methods, both of which are reliant on the well-known Bellman Equation, i.e.,
${\bm{V}}^{\pi}={\bm{r}}^{\pi}+\gamma{\bm{P}}^{\pi}{\bm{V}}^{\pi}.$ (2)
The first stream of methods, known as model-based estimation, is detailed in Algorithm 1. Model-based estimation $\widehat{\bm{V}}^{\textnormal{MB}}$ utilizes the reformulation of the Bellman Equation as ${\bm{V}}^{\pi}=({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-1}{\bm{r}}^{\pi}$, where ${\bm{I}}$ is the identity matrix. This method estimates ${\bm{V}}^{\pi}$ by plugging in the estimations of the MDP parameters including ${\bm{r}}^{\pi}$ and ${\bm{P}}^{\pi}$. Such an estimation is based on the comprehensive understanding of all model parameters, hence it is termed model-based.
The second methodology, referred to as model-free estimation, is outlined in Algorithm 2. It directly estimates the value function without estimating the model’s parameters. Specifically, for a given tuple $(s_{i},a_{i},s_{i+1},r_{i})$, $r_{i}+\gamma V(s_{i+1})$ is an unbiased estimation of $V(s_{i})$ according to the Bellman Equation. Thus, we obtain the estimator $\widehat{\bm{V}}^{\textnormal{TD}}$ following the well-known least square temporal difference (TD) algorithm, which solves
$\min_{\widehat{\bm{V}}}\sum_{s\in{\mathcal{S}}}\left({\sum_{s_{i}=s}r(s_{i},a_{i})+\gamma\widehat{V}(s_{i+1})-\widehat{V}(s_{i})}\right)^{2}.$ (3)
Despite of the difference in nature, $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ are actually providing the same point estimation of the value function ${\bm{V}}$. More formally, we have the following proposition.
Proposition 3.1
Let $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ be the respective output of Algorithm 1 and Algorithm 2 given the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ sampled under policy $\pi$. Assume all states are visited at least once in $\tau$, then it holds $\widehat{\bm{V}}^{\textnormal{MB}}=\widehat{\bm{V}}^{\textnormal{TD}}$.
Proof 3.2
Proof of Proposition 3.1. Let $N(s^{i})>0$ be the number of visitations to state $s^{i}\in{\mathcal{S}}$ in $\tau$. Then the optimization problem in Algorithm 2 can be rewritten as
$\min_{\widehat{\bm{V}}}\left\|{\operatorname{Diag}(N(s^{1}),N(s^{2}),\ldots,N(s^{K}))\cdot\left({\widehat{\bm{V}}-\left({\widehat{\bm{r}}+\gamma\widehat{\bm{P}}\widehat{\bm{V}}}\right)}\right)}\right\|_{2}^{2}\,,$ (4)
where $\widehat{\bm{r}}\in\mathbb{R}^{K}$ is the empirical average reward with $\widehat{\bm{r}}(s)=\sum_{i=1}^{T}\mathbf{1}[s_{i}=s]r(s_{i},a_{i})/N(s)$ for $s\in{\mathcal{S}}$ and $\widehat{\bm{P}}\in\mathbb{R}^{K\times K}$ is the empirical transition matrix with $\widehat{\bm{P}}(s,s^{\prime})=\sum_{i=1}^{T}\mathbf{1}[s_{i}=s,s_{i+1}=s^{\prime}]/N(s)$ for all $s,s^{\prime}\in{\mathcal{S}}$.
It is worth noting that any estimation $\widehat{\bm{P}}$ is always a transition matrix. Thus the matrix $({\bm{I}}-\gamma\widehat{\bm{P}})$ is always non-singular. Therefore one can verify that the unique solution to Eq. (4) is
$\widehat{\bm{V}}^{\textnormal{TD}}=({\bm{I}}-\gamma\widehat{\bm{P}})^{-1}\widehat{\bm{r}}\,,$
[…]
We have now demonstrated the equivalence between the two estimators within finite-space and finite action MDPs without function approximations, indicating that they share identical theoretical properties. Consequently, establishing a property for one estimator inherently confirms it for the other. It is important to emphasize, however, that this equivalence does not imply that the model-based approach and TD learning are identical across all scenarios. There has been a long debate in the reinforcement learning literature between model-based learning and TD learning, and we refer to the more relevant surveys and textbooks like Moerland et al. (2023) and Sutton and Barto (2018).
0:  The obtained trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$; Discount factor $\gamma$.
1:  $N(s,a)\leftarrow\sum_{i=1}^{T}\mathbb{I}\{s_{i}=s,a_{i}=a\}$, $N(s,a,s^{\prime})\leftarrow\sum_{i=1}^{T}\mathbb{I}\{s_{i}=s,a_{i}=a,s_{i+1}=s^{\prime}\}$, $\forall s,s^{\prime}\in{\mathcal{S}},a\in{\mathcal{A}}$
2:  $\widehat{r}(s)\leftarrow(\sum_{i=1}^{T}r_{i}\mathbb{I}\{s_{i}=s\})/(N(s)\vee 1)$, $\forall s\in{\mathcal{S}}$
13:  $\widehat{P}(s^{\prime}\mid s,a)=N(s,a,s^{\prime})/(N(s,a)\vee 1)$, $\forall s,s^{\prime}\in{\mathcal{S}},a\in{\mathcal{A}}$
4:  $\widehat{\bm{V}}^{\textnormal{MB}}\leftarrow({\bm{I}}-\gamma\widehat{\bm{P}})^{-1}\widehat{\bm{r}}$.
5:  return  $\widehat{\bm{V}}^{\textnormal{MB}}$.
Algorithm 1 Model-based Estimation
10:  The obtained trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$; Discount factor $\gamma$.
21:  Solve the following optimization problem Eq. (3),
$\min_{\widehat{\bm{V}}}\sum_{s\in{\mathcal{S}}}\left({\sum_{s_{i}=s}r(s_{i},a_{i})+\gamma\widehat{V}(s_{i+1})-\widehat{V}(s_{i})}\right)^{2}.$
2:  return  $\widehat{\bm{V}}^{\textnormal{TD}}$.
Algorithm 2 Model-free Estimation (Temporal Difference Learning)
3.1 Asymptotic Inference of Value Function
In this subsection, we investigate the asymptotic normality of the estimators $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$. For a fixed and static policy $\pi$, the MDP can induce a Markov chain where the states sequentially evolve. The states we observed on the trajectory $\tau$ are sampled from such a Markov chain, and the technical subtlety is that the samples are not necessarily independent and identically distributed (i.i.d.). Particularly, the current state $s_{t+1}$ depends on the last state $s_{t}$. However, if $s_{t}$ is sampled from the stationary distribution of the induced Markov chain, the future samples can be seen as i.i.d. generated from the stationary distribution. This observation allows us to hypothesize that rapid convergence to the stationary distribution may lead to asymptotic normality. To elucidate this concept further, we introduce the definition of uniform ergodicity, which indicates rapid mixing within the Markov chain. This principle is akin to those found in the works of Farias et al. (2022) and Hu and Wager (2022) and it is commonly assumed in the central limit theorem for Markov chains as detailed by Jones (2004).
Definition 3.3 (Uniform Ergodicity)
A Markov chain with the state transition matrix ${\bm{P}}$ is uniformly ergodic if it has stationary distribution ${\bm{\mu}}$ and there exists some constant $C\geq 0$ and $10\leq\lambda<$ 1 such that for all $k\in\mathbb{N}$ and all $s\in{\mathcal{S}}$,
$D_{\mathrm{TV}}({\bm{P}}^{k}(s,\cdot),{\bm{\mu}})\leq C\lambda^{k}\,,$
where $D_{\mathrm{TV}}(\cdot,\cdot)$ is the total variation distance.
Now we present the main theorem of this section, establishing the central limit theorem for $\widehat{\bm{V}}^{\textnormal{MB}}$, which naturally holds for $\widehat{\bm{V}}^{\textnormal{TD}}$. The detailed proof is delayed to Appendix 12.2.
Theorem 3.4 (Asymptotic Normality)
Assuming the Markov chain induced by policy $\pi$ is uniformly ergodic, with the input of a trajectory generated by a fixed policy $\pi$, it holds that as $T\to\infty$,
[…]
where
$\Sigma_{\mathrm{MB}}=({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-1}\operatorname{Diag}_{s\in{\mathcal{S}}}\left({\sigma^{2}_{\pi}(s)/\mu^{\pi}(s)+\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}/\mu^{\pi}(s)}\right)({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-\top}\,.$
We use $\sigma^{2}_{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}\mathbb{P}^{\pi}(a\mid s)\sigma^{2}(s,a)$ to denote the variance of the instant rewards of state $s$ under policy $\pi$. The vector ${\bm{\mu}}^{\pi}\coloneqq(\mu^{\pi}(s^{1}),\cdots,\mu^{\pi}(s^{K}))^{\top}$ is the stationary distribution under policy $\pi$, and $\Sigma_{p^{\pi}(s)}\coloneqq\operatorname{Diag}({\bm{p}}_{s})-{\bm{p}}_{s}{\bm{p}}_{s}^{\top}$, where ${\bm{p}}_{s}\in\mathbb{R}^{K}$ is the vector of the transition probability from state $s$ to the other states, $(\mathbb{P}^{\pi}(s^{1}\mid s),\cdots,\mathbb{P}^{\pi}(s^{K}\mid s))$.
Theorem 3.4 reveals that the model-based estimator is well-behaved and asymptotically normal. By Proposition 3.1, $\Sigma_{\mathrm{TD}}$ shares the same values and properties. In order to simplify and provide more insights, we discuss the inference of the value function under a specific state, say, the crucial state $s^{1}$. The following corollary presents the asymptotic normality of the estimator of $V(s^{1})$, which is the first dimension of ${\bm{V}}^{\textnormal{MB}}$ or ${\bm{V}}^{\textnormal{TD}}$.
Corollary 3.5
Following the same assumption as Theorem 3.4, it holds that as $T\to\infty$,
[…]
where
$\sigma^{2}_{\mathrm{MB}}(s^{1})=\sum_{s\in{\mathcal{S}}}\frac{\rho_{s^{1}}^{\pi}(s)^{2}}{\mu^{\pi}(s)}\left({\sigma_{\pi}^{2}(s)+\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}}\right)\,.$
Recall that $\rho^{\pi}_{s^{1}}(s)$ is the discounted state-visitation measure defined in Definition 2.1. The asymptotic variance of estimating $V(s^{1})$ is influenced not only by the state $s^{1}$ itself but also by all the other states. This is consistent with our understanding that our objective is the long-term value and the samples are drawn from a MDP. Although the variance depends on all states, we want to specially highlight that $\sigma^{2}_{\mathrm{MB}}(s^{1})$ can be decomposed into $K$ distinct terms, where each term corresponds almost exclusively to one state. Furthermore, each term is inversely proportion to the stationary probability $\mu^{\pi}(s)$. This implies that states which are less frequently visited pose greater challenges in sampling, potentially leading to increased variance. Additionally, each term has two parts, $\sigma_{\pi}^{2}(s)$ and $\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}$, which correspond to the (short-term) variance of the instant rewards and the (long-term) variance related to transitions to the other states.
3.2 Lower Bound for Asymptotic Variance of Estimating Value Function
In this subsection, we elaborate on the lower bound of the variance of any unbiased estimator. We reveal that the $\Sigma_{\mathrm{MB}}$ derived in Theorem 3.4 is optimal in the asymptotic view and cannot be further improved.
Theorem 3.6 (Asymptotic Efficiency)
Let the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ be sampled from MDP under policy $\pi$ with initial distribution ${\bm{\mu}}^{\pi}$. Then for any unbiased estimator of value functions with covariance matrix $\Sigma$, it holds that $\Sigma_{\mathrm{MB}}\preceq T\Sigma$, where $\Sigma_{\mathrm{MB}}$ is defined in Theorem 3.4.
To prove the theorem, let ${\bm{\theta}}=({\bm{P}}^{\pi},{\bm{r}}^{\pi})$ be the parameters controlling the system. Notice that ${\bm{\theta}}$ subjects to the constraint $\sum_{j}{\bm{P}}^{\pi}_{ij}=1$ for all $i\in[K]$. We then apply the Constrained Cramér–Rao bound (CCRB, see Lemma 18.5) for the lower bound. The full proof is delayed to Appendix 12.3.
In the recent literature, Khamaru et al. (2021) also establish a similar variance lower bound for TD learning, where they additionally assume the access to a generative model of the Markov chain. In particular, instead of sampling from a Markov chain, the generative model endows them with the flexibility to draw any random state transitions of the Markov chain. They can start from any state $s$ and generate the next state $s^{\prime}$ sampled from the distribution $\mathbb{P}^{\pi}(\cdot\mid s)$. With the generative model, they derive an asymptotic variance bound of TD as
$\Sigma_{\textnormal{GM}}=({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-1}\operatorname{Diag}_{s\in\mathcal{S}}\left({\sigma_{\pi}^{2}(s)+\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}}\right)({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-\top}\,.$
Different from our $\Sigma_{\textnormal{MB}}$, such a lower bound shaves off the stationary distribution ${\bm{\mu}}^{\pi}$ at the denominators. This matches the intuition that we can only sample from the Markov chain instead of from the generative model.
4 A/B Testing on MDPs
In the previous section, we have established the inference results with the trajectory generated by a fixed policy $\pi$. To infer the ATE, however, it is necessary to simultaneously estimate ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ from a single trajectory. A commonly employed method for this purpose is A/B testing, where, at any given state $s$, a fair coin flip determines whether to apply $\pi^{\mathfrak{t}}$ or $\pi^{\mathfrak{c}}$. For simplicity, we introduce the experimental policy $\pi^{1/2}$, defined such that $\pi^{1/2}(a\mid s)\coloneqq\frac{1}{2}\cdot\pi^{\mathfrak{t}}(a\mid s)+\frac{1}{2}\cdot\pi^{\mathfrak{c}}(a\mid s)$ for any $(s,a)$ pair. Assume that the experimenter executes $\pi^{1/2}$ for $T$ epochs and obtains the trajectory $\tau=\left\{{(s_{i},a_{i},s_{i+1},r_{i})}\right\}_{i\in[T]}$. An example of a realized trajectory might include sequences such as:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{3},{\mathfrak{c}},s^{3},r_{3})}\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
The blue and orange pairs are the observations generated by the treatment policy $\pi^{\mathfrak{t}}$ and the control policy $\pi^{\mathfrak{c}}$, respectively. Given that we are considering the SST, action ${\mathfrak{t}}$ is exclusively available at state $s^{1}$, while action ${\mathfrak{c}}$ is the only option in all other states, regardless of the policy applied. Without considering the SST structure—that is, treating $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ as two completely distinct and general policies—a straightforward approach would involve splitting the trajectory $\tau$ into two separate datasets based on the policy implemented at each time step:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}}=\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{t}},i\in[T]\}\,,\qquad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}}=\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{c}},i\in[T]\}\,,$
where $\pi_{i}$ denotes policy used at time $i$. Leveraging the Markov property of the system, unbiased estimations of ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ can be derived from ${\mathcal{D}}^{\mathfrak{t}}$ and ${\mathcal{D}}^{\mathfrak{c}}$ respectively, using either Algorithm 1 or Algorithm 2. The corresponding algorithm is outlined in Algorithm 3.
0:  Policy $\pi^{\mathfrak{t}},\pi^{\mathfrak{c}}$; Experiment horizon $T$; Discount factor $\gamma$.
1:  Execute the experiment policy $\pi^{1/2}$ for $T$ epochs and obtain datasets ${\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}=\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{\mathfrak{t}},i\in[T]\}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}=\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{\mathfrak{c}},i\in[T]\}$.
2:  Call Algorithm 1 or Algorithm 2 for $\widehat{{\bm{V}}^{\mathfrak{t}}}$ and $\widehat{{\bm{V}}^{\mathfrak{c}}}$ with ${\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}$ respectively.
3:  return  $\widehat{{\bm{V}}^{\mathfrak{t}}}-\widehat{{\bm{V}}^{\mathfrak{c}}}\,.$
Algorithm 3 A/B Testing on Markov Decision Process
In the previous section, we have demonstrated that both Algorithm 1 and Algorithm 2 exhibit asymptotic normality. Given that these algorithms are directly employed in Algorithm 3, it should not be surprising to anticipate that A/B testing within the MDP framework may also achieve asymptotic normality. Formally, we have the following theorem.
Theorem 4.1 (Asymptotic Normality of A/B Testing)
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic with stationary distribution ${\bm{\mu}}_{1/2}\coloneqq(\mu_{1/2}(s^{1}),\cdots,\mu_{1/2}(s^{K}))^{\top}$, it holds that as $T\to\infty$,
[…]
where
$\Sigma_{\mathrm{AB}}$ $=2({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-1}\operatorname{Diag}_{s\in{\mathcal{S}}}\left({\frac{\sigma_{\mathfrak{t}}^{2}(s)}{\mu_{1/2}(s)}+\frac{\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}}{\mu_{1/2}(s)}}\right)({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-\top}$
$\qquad+2({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{c}})^{-1}\operatorname{Diag}_{s\in{\mathcal{S}}}\left({\frac{\sigma_{\mathfrak{c}}^{2}(s)}{\mu_{1/2}(s)}+\frac{\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}{\mu_{1/2}(s)}}\right)({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{c}})^{-\top}\,.$
Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
[…]
where
$\sigma_{\mathrm{AB}}^{2}(s^{1})=2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}\Big{)}+2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)\,.$
Although similar to Theorem 3.4 and Corollary 3.5, the terms in $\Sigma_{\mathrm{AB}}$ and $\sigma_{\mathrm{AB}}^{2}(s^{1})$ in A/B testing are split into two distinct components: one associated exclusively with $\pi^{\mathfrak{t}}$ and the other with $\pi^{\mathfrak{c}}$. Each term in $\Sigma_{\mathrm{AB}}$ and $\sigma_{\mathrm{AB}}^{2}(s^{1})$ includes an additional constant factor of 2 compared to $\Sigma_{\mathrm{MB}}$ and $\sigma_{\mathrm{MB}}^{2}(s^{1})$ outlined in Theorem 3.4 and Corollary 3.5. This reflects the fact that A/B testing employs a mixed policy, with $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ each applied half of the time in expectation. Furthermore, the denominator features $\mu_{1/2}(s)$ rather than $\mu^{\pi^{\mathfrak{t}}}(s)$ or $\mu^{\pi^{\mathfrak{c}}}(s)$, indicating that the stationary distribution is linked to the mixed policy $\pi^{1/2}$.
Finally, we want to remind that Theorem 4.1 is applicable for any general $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$. Our approach segments the trajectory solely based on whether the policy $\pi^{\mathfrak{t}}$ or $\pi^{\mathfrak{c}}$ is used at each step, independent of what exactly $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ are. Consequently, this estimator can be readily extended to estimate ATE for a broader range of treatment policies $\pi^{\mathfrak{t}}$, beyond those confined to SST.
4.1 Lower Bound for Experimentation with General Treatment
In this subsection, we demonstrate that if we ignore the structure of SST, or if the treatment does not adhere to SST principles at all, then the asymptotic variance $\Sigma_{\textrm{AB}}$ given in Theorem 4.1 is almost optimal for any unbiased estimator of the ATE ${\bm{\Delta}}$. The proof is delayed to Appendix 13.2.
Theorem 4.2 (Asymptotic Efficiency of A/B Testing)
If running the mixed policy $\pi^{1/2}$ with the initial state distribution ${\bm{\mu}}_{1/2}$, for any unbiased estimator of ATE ${\bm{\Delta}}$ with covariance matrix $\Sigma$, it holds that $\Sigma_{\mathrm{AB}}\preceq T\Sigma$, where $\Sigma_{\mathrm{AB}}$ is defined in Theorem 4.1.
This lower bound effectively implies that asymptotic variance of A/B testing cannot be further reduced under general treatment, which is of independent interest to those who are considering experiments on discounted reward MDPs. In recent literature, Farias et al. (2022) derive a lower bound for A/B testing with general treatment on average-reward MDPs. However, their approach differs from ours in several key aspects. Firstly, their definition of ATE contrasts with ours; they define it as the difference between average rewards for $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ under stationary distribution, i.e. ${\bm{\mu}}^{{\mathfrak{t}}^{\top}}{\bm{r}}^{\mathfrak{t}}-{\bm{\mu}}^{{\mathfrak{c}}^{\top}}{\bm{r}}^{\mathfrak{c}}$. Additionally, our study further incorporates reward variance, whereas they assume that the reward observation is noise-free. From a technical perspective, our approach employs a distinct proof framework. We apply a constrained Cramér-Rao bound to establish the lower bound of variance for the transition matrix estimation. This represents a departure from the methodology used in Farias et al. (2022).
5 Variance Reduction: Information Sharing between Test Arms
In this section, we explore how the structure of SST can be leveraged to reduce variance. The underlying principle is that, thanks to the Markov property, the state transitions and reward signals of states $s_{t}\neq s^{1}$ can be shared across different test arms. To illustrate this, let us revisit the trajectory in Section 4. Although $(s^{3},{\mathfrak{c}},s^{3},r_{3})$ is generated by $\pi^{\mathfrak{c}}$, it can also be seen as having been generated by $\pi^{\mathfrak{t}}$ since $\pi^{\mathfrak{t}}(s^{3})=\pi^{\mathfrak{c}}(s^{3})={\mathfrak{c}}$. This observation suggests a potential strategy to use such samples for estimating both ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{t}}$. Consequently, we recolor the trajectory with additional purple on the samples that can be shared between the two policies.
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{3},{\mathfrak{c}},s^{3},r_{3})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
Accordingly, we can construct two datasets,
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}}=\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{t}})\}\cup\{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\},$
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}}=\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{c}})\}\cup\{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\}.$
Applying either Algorithm 1 or Algorithm 2 on the datasets ${\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}$ yields two new estimators $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}$ and $\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$ respectively. Subsequently, the estimator for ATE, denoted as $\widehat{\bm{\Delta}}^{\mathrm{IS}}$, is calculated as $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}-\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$. We will always use “IS” to denote information sharing. Observing any given trajectory $\tau$, it is evident that ${\mathcal{D}}^{\mathfrak{t}}_{\textrm{AB}}\subseteq{\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\textrm{AB}}\subseteq{\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}$. This suggests that information sharing is beneficial as it provides a larger dataset for fitting both ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ compared to the general A/B testing in Algorithm 3. However, one caveat here is that $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}$ and $\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$ may be sharing a significant portion of the data, introducing correlations between the estimations.
To assess whether information sharing can effectively reduce the variance, we first derive the central limit theorem with the information sharing. The full proof is delayed to Appendix 14.1.
Theorem 5.1 (Asymptotic Normality of Information Sharing)
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic, with SST, it holds that as $T\to\infty$,
[…]
where $\Sigma_{\mathrm{IS}}$ is provided in Appendix 14.1. Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
[…]
where
$\sigma_{\mathrm{IS}}^{2}(s^{1})=2$ $\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)$
$\qquad+\sum_{s\neq s^{1}}\frac{1}{\mu_{1/2}(s)}\left({\big{(}\rho^{\mathfrak{t}}_{s^{1}}(s)-\rho^{\mathfrak{c}}_{s^{1}}(s)\big{)}^{2}\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{\mathfrak{t}}-\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)\,.$
5.1 Analysis of Variance Reduction with Two Test Arms
In this subsection, we dive into the comparison between the asymptotic variance of general A/B testing derived in Theorem 4.1 and that of A/B testing with information sharing in Theorem 5.1. For simplicity, we restrict our focus to the crucial state $s^{1}$. Decomposing the variance based on whether the state is the crucial state or not, we can reorganize the asymptotic variance of A/B testing in Theorem 4.1 as
$\sigma_{\mathrm{AB}}^{2}(s^{1})$ $=2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}\Big{)}+2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)$
$=\underbrace{2\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)}_{\Sigma^{\mathrm{het}}_{\mathrm{AB}}}$
$\qquad+\underbrace{2\sum_{s\neq s^{1}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}\Big{)}+2\sum_{s\neq s^{1}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)}_{\Sigma^{\mathrm{hom}}_{\mathrm{AB}}}\,.$
The $\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ part is the variance directly related to the crucial state $s^{1}$, and the superscript het reflects that state $s^{1}$ is the only state where $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ may be heterogeneous under SST. The $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ part represents the variance associated with all other states $s\neq s^{1}$ and $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ are homogeneous. We want to remind that under SST, the variances $\sigma^{2}_{\mathfrak{t}}(s)$ and $\sigma^{2}_{\mathfrak{c}}(s)$ are the same for all the other state $s$ except the crucial state $s^{1}$.
Similarly, we can decompose the asymptotic variance of information sharing, $\sigma_{\mathrm{IS}}^{2}(s^{1})$, in the same fashion,
$\sigma_{\mathrm{IS}}^{2}(s^{1})$ $=\underbrace{2\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)}_{\Sigma^{\mathrm{het}}_{\mathrm{IS}}}$
$\qquad+\underbrace{\sum_{s\neq s^{1}}\frac{1}{\mu_{1/2}(s)}\left({\left({\rho^{\mathfrak{t}}_{s^{1}}(s)-\rho^{\mathfrak{c}}_{s^{1}}(s)}\right)^{2}\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{\mathfrak{t}}-\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)}_{\Sigma^{\mathrm{hom}}_{\mathrm{IS}}}\,.$
The above decomposition demonstrates clear structures of asymptotic variance of both estimators. Formally, the following Theorem 5.2 compares $\sigma_{\mathrm{AB}}^{2}(s^{1})$ and $\sigma_{\mathrm{IS}}^{2}(s^{1})$.
Theorem 5.2
Consider SST with treatment policy $\pi^{\mathfrak{t}}$ and control policy $\pi^{\mathfrak{c}}$, it holds
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{ and }\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}.$ (5)
Theorem 5.2 reveals that IS achieves at least an equivalent level of performance to the general A/B testing in terms of asymptotic variance. It aligns with our intuition that the variances associated with $s^{1}$ remain unchanged, as information sharing is not feasible in state $s^{1}$ due to the potential differences in the transition matrix and rewards. Consequently, the observed reduction in variance primarily results from the sharing of information across all other states $s\neq s^{1}$. We would like to highlight that, interestingly, each term of the sum in $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ can be directly related to each term of the sum in $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ through two Cauthy-Schwarz inequalities. More specifically, for any $s\neq s^{1}$,
$\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{\mathfrak{t}}-\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}$ $\leq 2\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{t}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}+2\left\|{\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{c}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2},$
$(\rho^{\mathfrak{t}}_{s^{1}}(s)-\rho^{\mathfrak{c}}_{s^{1}}(s))^{2}\sigma_{\mathfrak{c}}^{2}(s)$ $\leq 2\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}\sigma_{\mathfrak{t}}^{2}(s)+2\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}\sigma_{\mathfrak{c}}^{2}(s),$
where the sum of LHS over $s\neq s^{1}$ constitutes the $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ and the sum of RHS forms $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. Such a claim leading to the inequality in Eq. (5). To achieve equality in $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$, for all $s\neq s^{1}$, the two Cauthy-Schwarz inequalities above must be equal, which seems to be highly restrictive. In fact, Proposition 5.3 presents the necessary condition for the equality to hold.
Proposition 5.3 (Necessary Condition of Equality)
Equality in the inequality in Eq. (5) is achieved only if the following two conditions are satisfied simultaneously:
  1. 1.
    the rewards are fixed constants, i.e, $\sigma^{2}_{\mathfrak{c}}(s)=0$ for all $s\neq s^{1}$;
  2. 2.
    Either $r(s,a)=0$ for all $s\in{\mathcal{S}},a\in{\mathcal{A}}$ or $\exists\kappa\neq 0$ such that $\rho^{\mathfrak{t}}_{s^{1}}(s)=\kappa\rho^{\mathfrak{c}}_{s^{1}}(s)$ for all $s\neq s^{1}$.
The proof is delayed to Appendix 14.2. This corollary effectively implies that for the equality to hold: either the MDP is trivial with fixed zero reward at every state, or the discounted state visitation measure has to satisfy linear dependence before and after the treatment.
Additionally, we want to highlight that if the new treatment ${\mathfrak{t}}$ is identical to the control ${\mathfrak{c}}$ such that for all states $s$, $\rho^{\mathfrak{t}}_{s^{1}}(s)=\rho^{\mathfrak{c}}_{s^{1}}(s)$ and ${\bm{V}}^{\mathfrak{t}}={\bm{V}}^{\mathfrak{c}}$, the term $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ will be zero. In contrast, $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ is a value that can be strictly large than zero. This finding is particularly relevant in contexts like A/A testing. Such an observation is somewhat surprising. With information sharing, the amount of data used to estimate the transition probabilities and the rewards can be thought to be doubled, and thus it seems reasonable to conjecture that IS can at most reduce $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ to a half, i.e., $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\geq\frac{1}{2}\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. However, with the example of A/A testing, the conjecture may not always hold true, suggesting that $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ could sometimes be significantly smaller than $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$.
5.2 Lower Bound for Experimentation with SST
In this subsection, we present the variance lower bound of any unbiased estimator taking the SST structure into account, which reflects the intrinsic difficulty of conducting inference with SST.
Theorem 5.4 (Asymptotic Efficiency of Information-Sharing)
Assume the treatment is SST with two test arms. If running mixed policy $\pi^{1/2}$ with initial distribution $\mu_{1/2}$, for any unbiased estimator of ATE ${\bm{\Delta}}$ with covariance matrix $\Sigma$, it holds $\Sigma_{\mathrm{IS}}\preceq T\Sigma$, where $\Sigma_{\mathrm{IS}}$ is defined in Theorem 5.2.
Theorem 5.4 indicates that our estimator $\widehat{\bm{\Delta}}^{\mathrm{IS}}$ is asymptotically efficient. While Theorem 5.2 highlights that IS does not reduce the variance associated with the crucial state $s^{1}$, Theorem 5.4 affirms that this variance is, in fact, optimal and fundamental.
5.3 (More) Power of IS with Multiple Test Arms
In previous sections, we have shown that IS can do at least as well as A/B testing with two test arms. A practical experiment often contains more than two test arms, such as A/B/n Testing (Kohavi et al. 2020, Russac et al. 2021) or bandit experiments (Simchi-Levi and Wang 2023). In this subsection, we generalize our study of the variance reduction of IS to SST with more than two test arms and show that IS becomes even more powerful in reducing the variance comparing with traditional A/B/n testing. For rigor, we start with the definition of SST with $n$ test arms.
Consider a MDP with state space ${\mathcal{S}}\coloneqq\left\{{s^{i}}\right\}_{i\in[K]}$, where $s^{1}$ is still assumed to be the crucial state of treatment. Unlike the previous model in Section 2 with two test arms, the action set ${\mathcal{A}}=\left\{{{\mathfrak{t}}^{j}}\right\}_{j\in[n]}$ now includes $n$ possible actions. Without loss of generality, we let ${\mathfrak{t}}^{1}$ denote the control action. Recall Example 1.1, a plausible experiment might involve the company testing multiple coupon values to determine which one maximizes customer lifetime value. This implies $n$ SST polices $\Pi\coloneqq\left\{{\pi^{i}}\right\}_{i\in[n]}$ with each $\pi^{i}$ satisfying: $\pi^{i}(s^{1})\equiv{\mathfrak{t}}^{i}$ (i.e., $\pi^{i}({\mathfrak{t}}^{i}\mid s^{1})=1$) and $\pi^{i}(s)\equiv{\mathfrak{t}}^{1}$ (i.e., $\pi^{i}({\mathfrak{t}}^{1}\mid s)=1$) for all $s\in{\mathcal{S}}\setminus s^{1}$. Additionally, each SST policy is associated with its own transition matrix ${\bm{P}}^{{\mathfrak{t}}^{i}}$ and mean reward ${\bm{r}}^{{\mathfrak{t}}^{i}}$.
The primary objective of experimentation in such a setting is to understand the long-term outcome differences among various test arms. Formally, for any pair of SST polices $(\pi^{i},\pi^{j})_{i,j\in[n]}$, we define the ATE as $\Delta_{i,j}(s)=V^{i}(s)-V^{j}(s)$, for given state $s\in{\mathcal{S}}$. For simplicity, throughout this subsection, we restrict our interest to $\Delta_{i,j}(s^{1})$, the ATE under the crucial state $s^{1}$. For A/B/n testing, the experimental policy becomes $\pi^{1/n}(a\mid s)\coloneqq\frac{1}{n}\sum_{i\in[n]}\pi^{i}(a\mid s)$ for any $(s,a)$ pair, meaning that we apply each policy in $\Pi$ with equal probability.
After collecting data, the traditional A/B/n testing would split the trajectory according to different polices and obtain dataset ${\mathcal{D}}^{j}_{\mathrm{ABn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{j},i\in[T]\}$ for each policy $\pi^{j}$. With information sharing, the observations with $s_{i}\neq s^{1}$ among all $n$ test arms can be shared for all SST polices $\pi$. Then, IS yields the corresponding dataset ${\mathcal{D}}^{j}_{\mathrm{ISn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):(s_{i},a_{i})\neq(s^{1},{\mathfrak{t}}^{k}),\forall k\neq j,i\in[T]\}$ for each policy $\pi^{j}$. Applying Algorithm 1 or Algorithm 2 to fit ${\bm{V}}^{j}$ with ${\mathcal{D}}^{j}_{\mathrm{ABn}}$ and ${\mathcal{D}}^{j}_{\mathrm{ISn}}$ respectively can produce two estimators $\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The estimators for the treatment effect become $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ABn}}=\widehat{\bm{V}}^{\textnormal{ABn}}_{i}-\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ISn}}=\widehat{\bm{V}}^{\textnormal{ISn}}_{i}-\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The following Corollary 5.5 generalizes Theorems 4.1 and 5.1 revealing the asymptotic normality of these two estimators on the crucial state $s^{1}$.
Corollary 5.5 (Asymptotic Normality of A/B/n Testing with(out) IS)
Assuming the Markov Chain induced by the mixed policy $\pi^{1/n}$ is uniformly ergodic, with SST, it holds as $T\to\infty$, for both $\Gamma\in\{\mathrm{ABn},\mathrm{ISn}\}$,
[…]
where
$\sigma_{\mathrm{ABn},(i,j)}^{2}(s^{1})$ $=\underbrace{n\left({\frac{\rho^{i}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{i}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{i}}\right\|_{\Sigma_{p^{i}}(s^{1})}^{2}}\right)+\frac{\rho^{j}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{j}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{j}}\right\|_{\Sigma_{p^{j}}(s^{1})}^{2}}\right)}\right)}_{\Sigma_{\mathrm{ABn}}^{\mathrm{het}}(i,j)}$
$\qquad+\underbrace{n\sum_{s\neq s^{1}}\frac{\rho^{i}_{s^{1}}(s)^{2}}{\mu_{1/n}(s)}\left({\sigma_{i}^{2}(s)+\left\|{\gamma{\bm{V}}^{i}}\right\|_{\Sigma_{p^{i}}(s)}^{2}}\right)+n\sum_{s\neq s^{1}}\frac{\rho^{j}_{s^{1}}(s)^{2}}{\mu_{1/n}(s)}\left({\sigma_{j}^{2}(s)+\left\|{\gamma{\bm{V}}^{j}}\right\|_{\Sigma_{p^{j}}(s)}^{2}}\right)}_{\Sigma_{\mathrm{ABn}}^{\mathrm{hom}}(i,j)},$
and
$\sigma_{\mathrm{ISn},(i,j)}^{2}(s^{1})=$ $\underbrace{n\left({\frac{\rho^{i}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{i}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{i}}\right\|_{\Sigma_{p^{i}}(s^{1})}^{2}}\right)+\frac{\rho^{j}_{s^{1}}(s^{1})^{2}}{\mu_{1/n}(s^{1})}\left({\sigma_{j}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{j}}\right\|_{\Sigma_{p^{j}}(s^{1})}^{2}}\right)}\right)}_{\Sigma_{\mathrm{ISn}}^{\mathrm{het}}(i,j)}$
$\qquad+\underbrace{\sum_{s\neq s^{1}}\frac{1}{\mu_{1/n}(s)}\left({\big{(}\rho^{i}_{s^{1}}(s)-\rho^{j}_{s^{1}}(s)\big{)}^{2}\sigma_{1}^{2}(s)+\left\|{\gamma\rho^{i}_{s^{1}}(s){\bm{V}}^{i}-\gamma\rho^{j}_{s^{1}}(s){\bm{V}}^{j}}\right\|_{\Sigma_{p^{1}}(s)}^{2}}\right)}_{\Sigma_{\mathrm{ISn}}^{\mathrm{hom}}(i,j)}\,.$
We ensure that the both two estimators $\widehat{\Delta}_{i,j}^{\mathrm{ABn}}(s^{1})$ and $\widehat{\Delta}_{i,j}^{\mathrm{ISn}}(s^{1})$ enjoy the asymptotic normality, but with different asymptotic variances. Similar as before, the variance can be decomposed according to states. Theorem 5.6 below characterizes the variance reduction of IS for multiple test arms demonstrating the more power that IS has when facing multiple test arms.
Theorem 5.6
Consider SST with $n$ test arms $\left\{{\pi^{i}}\right\}_{i\in[n]}$, it holds for any pair of test arms $(\pi^{i},\pi^{j})$,
$\Sigma^{\mathrm{het}}_{\mathrm{ISn}}(i,j)=\Sigma^{\mathrm{het}}_{\mathrm{ABn}}(i,j)\,,\qquad\Sigma^{\mathrm{hom}}_{\mathrm{ISn}}(i,j)\leq\frac{2}{n}\Sigma^{\mathrm{hom}}_{\mathrm{ABn}}(i,j),$ (6)
where the equality of the second part can hold only if the reward of states $s\in{\mathcal{S}}\setminus s^{1}$ are fixed, i.e. $\sigma_{{\mathfrak{t}}^{1}}^{2}(s)=0$ for all $s\neq s^{1}$.
This result underscores the effectiveness of IS in leveraging the structure of SST to significantly reduce the variance compared to traditional A/B/n testing. Particularly, we demonstrate that the variance associated with all other states $s\neq s^{1}$, where the control policy is always employed, can be reduced at least linearly with the number of test arms. Thus, the more test arms we have, the greater the reduction in variance achieved by IS. When $n=2$, Theorem 5.6 replicates the findings of Theorem 5.2. When $n\geq 3$, Theorem 5.6 consistently ensures a substantially reduced variance. Remarkably, with $n\geq 4$, IS is expected to cut the variance associated with the non-crucial states by at least half. Similar to Theorem 5.2, the variance associated with crucial state $s^{1}$ is the same as A/B/n testing, as information about this state cannot be shared. Additionally, the equality condition in the inequality of Eq. (6) remains as elusive as before, largely because one necessary condition for equality is the absence of noise in the observed rewards.
6 Estimation with Perfect Information about the Control Arm
In previous sections, we demonstrated how sharing information among test arms can effectively reduce variance. In this section, we explore this concept further by considering scenarios where we have Perfect Information (PI) about the control arm. Typically, the control arm represents a default policy that may have been operational for an extended period within the system, allowing the organization to accumulate a substantial dataset and develop a thorough understanding of this baseline. Extending Example 1.1, the company might already possess ample historical data on customer behaviors, enabling it to accurately determine the customers’ transition matrices when not offering coupon. We assume the historical data accurately characterizes the customer’s behavior (the same transition) at present. In other words, there is no distribution shift between the offline dataset and the current trial. Access to extensive historical user data is a common strategy in marketing, utilized to optimize selling policies (Simester et al. 2006, Yang et al. 2024), predict future purchases (Jacobs et al. 2016), or gauge customer responses (Narayandas and Sengupta 2023). Similarly, the assumption of having an offline dataset is prevalent in other domains involving MDPs, such as offline reinforcement learning (Levine et al. 2020), reinforcement learning with human feedback (RLHF, Chang et al. 2024), and healthcare applications (Yu et al. 2021, Hu et al. 2024).
Formally, we assume that the experimenter has access to both the precise transition matrix ${\bm{P}}^{\mathfrak{c}}$ and the exact mean reward ${\bm{r}}^{\mathfrak{c}}$ of control policy $\pi^{\mathfrak{c}}$. With this perfect information, she can directly computing ${\bm{V}}^{\mathfrak{c}}$ using the Bellman equation. Consequently, her primary focus during experimental periods becomes gathering data on the new treatment policy $\pi^{\mathfrak{t}}$, applying it consistently. Moreover, since $\pi^{\mathfrak{t}}$ is SST, a substantial amount of information from ${\bm{r}}^{\mathfrak{c}}$ and ${\bm{P}}^{\mathfrak{c}}$, especially those associated with the non-crucial states, can be shared when estimating ${\bm{V}}^{\mathfrak{t}}$. More specifically, after consistently applying $\pi^{\mathfrak{t}}$, she only needs to estimate the transitions and mean reward associated with the crucial state $s^{1}$, which is
$\widehat{{\bm{P}}}^{{\mathfrak{t}}}_{\textrm{PI}}(s^{1},s)=\frac{\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1},s_{i+1}=s]}{1\vee\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1}]}\,\forall s\in\mathcal{S}\quad\textrm{ and }\quad\widehat{r}_{\textrm{PI}}^{\mathfrak{t}}(s^{1})=\frac{\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1}]r_{i}}{1\vee\sum_{i=1}^{T}\mathbf{1}[s_{i}=s^{1}]}.$
For all the other components in $\widehat{{\bm{P}}}_{\textrm{PI}}^{{\mathfrak{t}}}$ and $\widehat{{\bm{r}}}_{\textrm{PI}}^{{\mathfrak{t}}}$, we directly plug in those in ${\bm{P}}^{\mathfrak{c}}$ and ${\bm{r}}^{\mathfrak{c}}$, i.e., $\widehat{{\bm{P}}}^{{\mathfrak{t}}}_{\textrm{PI}}(s,s^{\prime})={\bm{P}}^{\mathfrak{c}}(s,s^{\prime})$ and $\widehat{r}_{\textrm{PI}}^{\mathfrak{t}}(s)=r^{\mathfrak{c}}(s)$ for all $s\neq s^{1}$ and any $s^{\prime}\in\mathcal{S}$. Applying $\widehat{{\bm{P}}}_{\textrm{PI}}^{{\mathfrak{t}}}$ and $\widehat{{\bm{r}}}_{\textrm{PI}}^{{\mathfrak{t}}}$ to the Bellman equation generates an estimator $\widehat{{\bm{V}}}^{\mathfrak{t}}_{\textrm{PI}}$. Then, the difference between $\widehat{{\bm{V}}}^{\mathfrak{t}}_{\textrm{PI}}$ and ${{\bm{V}}}^{\mathfrak{c}}$ yields an estimator $\widehat{\bm{\Delta}}^{\mathrm{PI}}$ of the treatment effects. The following theorem presents the central limit theorem of our estimator. The detailed proof is delayed to Appendix 15.1.
Theorem 6.1 (Asymptotic Normality of Perfect Information)
Assuming the Markov chain induced by the SST policy $\pi^{\mathfrak{t}}$ is uniformly ergodic, it holds that as $T\to\infty$,
[…]
where
$\Sigma_{\mathrm{PI}}=({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-1}\begin{bmatrix}\frac{\sigma_{\mathfrak{t}}^{2}(s^{1})}{\mu^{\mathfrak{t}}(s^{1})}+\frac{\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}}{\mu^{\mathfrak{t}}(s^{1})}&\\ &{\bm{0}}\end{bmatrix}({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-\top}\,.$
Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
[…]
where
$\sigma^{2}_{\mathrm{PI}}(s^{1})=\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu^{\mathfrak{t}}(s^{1})}\left({\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|^{2}_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}}\right)\,.$
Since we are implementing a fixed policy $\pi^{\mathfrak{t}}$ during the experimental periods, it is logical that the asymptotic variance $\Sigma_{\mathrm{PI}}$ and $\sigma^{2}_{\mathrm{PI}}(s^{1})$ exhibit similar structures with $\Sigma_{\mathrm{MB}}$ and $\sigma^{2}_{\mathrm{MB}}(s^{1})$ from the model-based estimator under a fixed policy discussed in Section 3. Unlike the approach in that section, which requires estimating transitions for every state, here we only need to estimate the transitions and rewards associated with the crucial state $s^{1}$. Recalling Corollary 3.5, the asymptotic variance $\sigma^{2}_{\textrm{MB}}(s^{1})$ is the sum of variances associated with all states. With PI, $\sigma^{2}_{\textrm{PI}}(s^{1})$ only keeps the variance linked to the crucial state $s^{1}$, getting rid of all the other terms in $\sigma^{2}_{\textrm{MB}}(s^{1})$.
This reduction aligns with our expectations, given the complete information on the transitions and rewards for non-crucial states now at our disposal.
We now would like to compare the variance with A/B testing, both with and without information sharing, as discussed in Sections 4 and 5. In addition to the perfect information about the control arm, as we are running $\pi^{\mathfrak{t}}$ all the time instead of the mixed policy like A/B testing, we naturally accumulate more data specific to $\pi^{\mathfrak{t}}$. Hence it’s natural to anticipate a further reduction of variance associated with crucial state $s^{1}$. To facilitate a detailed comparison and ensure consistency in our analysis, we introduce two notations, albeit redundant, $\Sigma^{\mathrm{het}}_{\mathrm{PI}}$ and $\Sigma^{\mathrm{hom}}_{\mathrm{PI}}$. These notations are designed to differentiate the parts of $\sigma^{2}_{\mathrm{PI}}(s^{1})$ related to the crucial states $s^{1}$ and all other states, respectively. Thus, $\Sigma^{\mathrm{het}}_{\mathrm{PI}}=\sigma^{2}_{\mathrm{PI}}(s^{1})$ and $\Sigma^{\mathrm{hom}}_{\mathrm{PI}}=0$ since $\sigma^{2}_{\mathrm{PI}}(s^{1})$ is only related to $s^{1}$. Formally, we have the following theorem characterizing the variance reduction of PI, whose detailed proof is delayed to Appendix 15.2.
Theorem 6.2
Consider SST with treatment policy $\pi^{\mathfrak{t}}$ and control policy $\pi^{\mathfrak{c}}$, it holds
$\Sigma^{\mathrm{het}}_{\mathrm{PI}}<\left\{{1\land\frac{1+\delta L}{2}}\right\}\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{PI}}=0\leq\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}\,,$
where $\delta=D_{\mathrm{TV}}({\bm{P}}^{\mathfrak{t}}(s^{1},\cdot),{\bm{P}}^{\mathfrak{c}}(s^{1},\cdot))$ and $L\coloneqq(\ln(C)+1)/(1-\lambda)$. The constants $C,\lambda$ are defined in the uniform ergodicity of the Markov chain (Definition 3.3) induced by $\pi^{1/2}$.
The parameter $\delta$ measures the difference between the transition probabilities under the treatment and the control. One of the reasons that such a parameter takes an effect is that the denominators in $\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ and $\Sigma^{\mathrm{het}}_{\mathrm{PI}}$ are $\mu_{1/2}(s^{1})$ and $\mu^{{\mathfrak{t}}}(s^{1})$ respectively, which are the stationary probabilities of the crucial state $s^{1}$. A little difference in the transition probability may have an significant impact on the reciprocal of the stationary probability. Also, recall that $\Sigma^{\mathrm{het}}_{\mathrm{AB}}=\Sigma^{\mathrm{het}}_{\mathrm{IS}}$, and Theorem 6.2 confirms that PI can always strictly outperform A/B testing even with information sharing. Moreover, if the shift of transition probability $\delta$ is small or the Markov chain mixes fast enough, Theorem 6.2 indicates that PI can approximately halve the variance compared to A/B testing for the heterogeneous component. To reiterate, PI eliminates uncertainty stemming from all non-crucial states.
Similar ideas can also be generalized to SST with multiple test arms, where the perfect information about control arm is available. Following the notations in Section 5.3, assume there are $n$ test arm with the first arm serving as the control. Consequently, sampling is required only for the $n-1$ test arms excluding the control arm, such that $\pi^{1/(n-1)}(a\mid s)\coloneqq\frac{1}{n-1}\sum_{i\neq 1}\pi^{i}(a\mid s)$ for any $(s,a)$ pair. For all the test arms, we estimate only ${\bm{P}}^{{\mathfrak{t}}^{i}}(s^{1},\cdot)$ and $r^{{\mathfrak{t}}^{i}}(s^{1})$ using sample means, while inheriting the other values from the known control arm. Using the Bellman equation, we derive the estimator $\widehat{{\bm{V}}}^{{\mathfrak{t}}^{i}}$ for each SST, leading to the estimator of $\Delta_{i,j}(s^{1})$ by taking the difference. In Corollary 6.3, we establish the asymptotic normality for the estimation of $\Delta_{i,j}(s^{1})$, confirming the statistical robustness of our approach under these settings.
Corollary 6.3
Assuming the Markov chain induced by the mixed policy $\pi^{1/(n-1)},\pi^{1/n}$ are both uniformly ergodic, it holds that as $T\to\infty$,
[…]
where $\sigma_{\mathrm{PIn},(i,j)}^{2}(s^{1})$ is given in Appendix 15.1.1. Furthermore, we introduce $\Sigma^{\mathrm{het}}_{\mathrm{PIn}}(i,j)$ and $\Sigma^{\mathrm{hom}}_{\mathrm{PIn}}(i,j)$ to denote the terms in $\sigma_{\mathrm{PIn},(i,j)}^{2}(s^{1})$ associated with $s^{1}$ and all the other states respectively. We have
$\Sigma^{\mathrm{het}}_{\mathrm{PIn}}(i,j)<\left\{{1\land\frac{n-1}{n}\left({1+\frac{2\delta L}{n}}\right)}\right\}\Sigma^{\mathrm{het}}_{\mathrm{ABn}}(i,j)\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{PIn}}(i,j)=0\,,$
where $\delta=\max_{i}D_{\mathrm{TV}}({\bm{P}}^{i}(s^{1},\cdot),{\bm{P}}^{1}(s^{1},\cdot))$ and $L\coloneqq(\ln(C)+1)/(1-\lambda)$. The constants $C,\lambda$ are related to the uniform ergodicity of the Markov chain induced by $\pi^{1/n}$ from Definition 3.3.
With $n$ test arms, PI continues to yield asymptotically normal estimators and significant variance reduction. Moreover, the term $\left\{{1\land\frac{n-1}{n}\left({1+\frac{2\delta L}{n}}\right)}\right\}$ is almost non-decreasing with increasing $n$. This pattern indicates that as $n$ grows, the relative benefit of variance reduction for $\Sigma^{\mathrm{het}}_{\mathrm{PIn}}(i,j)$ diminishes. Such an observation aligns with the intuition that as the number of test arms grows, the data available for each test arm decreases, potentially becoming a limiting factor in further reducing variance.
7 From SST to Local Treatment
In practice, a local treatment policy may affect more than a single state. Consequently, we generalize the SST concept to accommodate Local Treatment. In this expanded framework, a treatment action is applied to a specific segment of the state space rather than a single state. This approach allows for a more flexible and realistic modeling of interventions in complex MDPs. To illustrate, recall Example 1.1. The company now wants to test a new strategy: posting advertisements to customers when their recency is relatively high—for instance, when it exceeds 2. Such a treatment policy is a Local Treatment, as it is applied only to a specific segment ($s^{3},s^{4},s^{5}$) of the state space. The remaining states ($s^{1},s^{2}$), representing lower recency, are not subject to this treatment.
We then establish the model for Local Treatment. Specifically, there exists a pre-determined “crucial” subset ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$. The treatment policy $\pi^{\mathfrak{t}}$ and control policy $\pi^{\mathfrak{c}}$ will only vary at state $s\in{\mathcal{S}}^{\mathfrak{t}}$. Formally, we define
  • The Control Policy $\pi^{\mathfrak{c}}$: This policy maintains $\pi^{\mathfrak{c}}(s)\equiv{\mathfrak{c}}$ across all states $s\in{\mathcal{S}}$, following the same definition as Section 2.
  • The Treatment Policy $\pi^{\mathfrak{t}}$: This policy now implements the new intervention ${\mathfrak{t}}$ exclusively whenever at the state $s\in{\mathcal{S}}^{\mathfrak{t}}$, while defaulting to ${\mathfrak{c}}$ in all the other states in ${\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$. In other words, $\pi^{\mathfrak{t}}(s)\equiv{\mathfrak{t}}$ for all $s\in{\mathcal{S}}^{\mathfrak{t}}$ and $\pi^{\mathfrak{t}}(s)\equiv{\mathfrak{c}}$ for all $s\in{\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$.
Notice that when the size of the crucial subset $|{\mathcal{S}}^{\mathfrak{t}}|=1$, we recover the definition of SST. The average treatment effect ${\bm{\Delta}}$ follows the same definition as section 2.
Having demonstrated the effectiveness of IS and PI for variance reduction under SST, we can now extend these results to Local Treatment scenarios. It can be expected that all A/B testing methods—including IS and PI—under local treatment will still produce asymptotically normal estimators, similar to Theorem 4.1, 5.1 and 6.1. Compared to SST, the heterogeneous part of the system becomes a subset ${\mathcal{S}}^{\mathfrak{t}}$ instead of the single just state $s^{1}$. As a result, we can only share information, or plug in perfect information, of state $s\in{\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$ for IS and PI estimator respectively. The following theorem characterizes this key difference from SST.
Theorem 7.1 (Asymptotic Normality with Local Treatment)
Assuming the Markov Chain induced by policy $\pi^{\mathfrak{t}},\pi^{1/2}$ are both uniformly ergodic, with Local Treatment on state ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$, it holds as $T\to\infty$, for $\Gamma\in\{\mathrm{AB},\mathrm{IS},\mathrm{PI}\}$ and any state $s\in{\mathcal{S}}$,
[…]
where $\sigma_{\Gamma}^{2}(s)$ can be decomposed according to the state as
$\sigma_{\Gamma}^{2}(s)$ $=\Sigma_{\Gamma}^{\mathrm{het}}+\Sigma_{\Gamma}^{\mathrm{hom}}\,,$
where $\Sigma_{\Gamma}^{\mathrm{het}}$ and $\Sigma_{\Gamma}^{\mathrm{hom}}$ refers to the variance of states ${\mathcal{S}}^{\mathfrak{t}}$ and ${\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$ respectively for $\Gamma\in\{\mathrm{AB},\mathrm{IS},\mathrm{PI}\}$. The full expression of $\sigma_{\mathrm{AB}}^{2}(s),\sigma_{\mathrm{IS}}^{2}(s),\sigma_{\mathrm{PI}}^{2}(s)$ can be found in Appendix 16.
Recall in SST, $\Sigma_{\Gamma}^{\mathrm{het}}$ is associated with the single crucial state $s^{1}$, while now we generalize it to the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$. For the asymptotic relative efficiency of three estimators, we then have the following generalization of Theorem 5.2 and Theorem 6.2.
Theorem 7.2
Consider Local Treatment Effect on ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$ with two test arms $\pi^{\mathfrak{t}},\pi^{\mathfrak{c}}$, it holds
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{PI}}=0\leq\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}\,.$
Theorem 7.2 replicates Theorem 5.2. The difference is that now we generalize $\Sigma_{\Gamma}^{\mathrm{het}}$ to variance associated with the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$ rather than single crucial state $s^{1}$ in SST. This demonstrates that the effectiveness of IS diminishes as the number of states that can share information (i.e. the homogeneous part) decreases. In the extreme case of a general treatment where ${\mathcal{S}}^{\mathfrak{t}}={\mathcal{S}}$, meaning all states are treated and no states to share information, IS will reduce to A/B testing. One caveat here is that unlike Theorem 6.2, we do not have general guarantees like $\Sigma^{\mathrm{het}}_{\mathrm{PI}}<\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ for local treatment. The intuition is that $\mu^{\mathfrak{t}}(s)$ can be much smaller than $\frac{1}{2}\mu_{1/2}(s)$ for some state $s$ when considering the local treatment. Hence this brings about large variance for the estimation of $\widehat{\Delta}^{\mathrm{PI}}(s)$. We leave it to future work for efficient design of experiments with perfect information and local treatment. Finally, we can derive a similar generalization to local treatment with multiple test arms as Theorem 5.6 and Corollary 6.3. We delay the details to the Appendix 16.1.
8 Simulations
We revisit the example of customer lifetime value (Example 1.1), with a synthetic transition matrix detailed in Appendix. We assume that under the control policy, when the customer makes a purchase, the company will receive a net income of $30$ while the income in all other states is 00. We consider the following two types of treatment:
  1. 1.
    SST. The company sends a coupon when the customer is at highest recency (recency=5).
  2. 2.
    Local Treatment. The company starts posting advertisements whenever the customer’s recency is high (recency$>$2).
For each type of treatment, we consider $5$ candidate coupon values (advertising cost), i.e. $n=5$ test arms (the coupon value (advertising cost) is 00 for the control arm). We then run experiments for each treatment type, aiming to identify the optimal coupon value (advertising cost) and report their ATE. In each experiment, we run all three estimators over 1,000 independent trajectories and record their empirical mean and standard error. The experiment horizon is adjusted so that the IS estimator’s standard error is smaller than its empirical mean.
Refer to caption Refer to caption
Figure 1: Synthetic simulations for Example 1.1. Diamonds and caps represent the empirical mean and standard error respectively for each estimator. Left: the result for SST policy with $n=5$ test arms and experiment horizon $T=350$. Right: the result for Local Treatment policy with $n=5$ test arms and experiment horizon $T=2000$.
Figure 1 summarizes the simulation results, we immediately see:
Unbiased Estimation Across All Estimators: The empirical means for all three estimators in both experiments demonstrate negligible bias. This aligns with our theoretical prediction of asymptotic normality for all three estimators.
IS and PI work: In both examples, we observe a significant reduction in variance—by at least half—through the application of IS and PI. This corroborates the effectiveness of IS and PI in overcoming the lower bound of A/B testing by utilizing the local structure. Moreover, PI consistently outperforms IS in both examples, confirming its capability in further variance reduction.
Local Treatment is Harder than SST: While ATEs of each experiment are close, the local treatment required a substantially longer horizon for the standard error of IS to decay compared to SST. This discrepancy arises because the SST allows information sharing for more states compared to local treatment. This comparison clearly illustrates that the efficacy of IS increases with the number of shared states.
9 Discussion
9.1 Fairness And Welfare: The Monotonicity of SST
MDPs are complex systems with interconnected states. Therefore, out of fairness and welfare concern, it would be the best for the decision-makers to ensure that a new treatment improves long-term outcomes for customers across all states, rather than benefiting some at the expense of others. Fortunately, we have a theorem that guarantees the “monotonicity” of SST policies. This ensures that the treatment effect remains consistently positive (or negative) for all states before and after the intervention.
Theorem 9.1 (Monotonicity of Single-State Treatment Effect)
Consider SST policy $\pi^{\mathfrak{t}}$ with crucial state $s^{1}$. If there exists $s\in{\mathcal{S}}$ such that $\Delta(s)>0$, then for all $s\neq s^{1}$,
$0\leq\Delta(s)<\gamma\Delta(s^{1})\,.$
Vice versa, if there exists $s\in{\mathcal{S}}$ such that $\Delta(s)<0$, then $\gamma\Delta(s^{1})<\Delta(s)\leq 0$ for all $s\neq s^{1}\,.$
The monotonicity of SST ensures that an improvement in any state will positively impact the long-term outcomes of all other states. Conversely, when the treatment effect is negative, all states except the crucial state will experience a welfare loss strictly less than $\gamma\Delta(s^{1})$, where $\gamma\in[0,1)$ is the discount factor. Additionally, the crucial state will always have the strongest treatment effect, aligning with our intuition that the crucial state is the primary target of the new treatment. Notably, this monotonicity theorem holds true regardless of the underlying Markov chain’s structure.
We provide two different proofs for Theorem 9.1 in Appendix 17. The first proof is based on an important observation that the treatment effect ${\bm{\Delta}}$ is an eigenvector of a special matrix associated with the transition probability under the control policy. By applying the Perron-Frobenius Theorem Meyer (2001) and carefully verifying the conditions, we arrive at the desired result. The second proof offers a novel perspective to characterize ${\bm{\Delta}}$ using the advantage function, derived from the well-knwon Performance Difference Lemma (Kakade and Langford 2002, Agarwal et al. 2022) in RL theory. We believe these techniques may be of independent interest to the RL literature.
9.2 Bias-Variance Trade-off
With general treatment, the seminal work by Farias et al. (2022) proposes a Difference-in-Q (DQ) estimator that trades a small bias for a striking variance reduction in average reward setting. In a follow-up work (Farias et al. 2023), they extend the idea of DQ to other reward settings including discounted reward used in this paper. In particular, we adapt their DQ theory to obtain the following estimator
$\widehat{\bm{\Delta}}^{\textrm{DQ}}\coloneqq\left({{\bm{I}}-\gamma\widehat{\bm{P}}_{1/2}}\right)^{-1}\left({\widehat{\bm{Q}}^{\pi_{1/2}}({\mathfrak{t}})-\widehat{\bm{Q}}^{\pi_{1/2}}({\mathfrak{c}})}\right)\,,$
where $\widehat{\bm{Q}}^{\pi_{1/2}}(a)=[\widehat{Q}^{\pi_{1/2}}(s^{i},a)]^{\top}_{i\in[K]}$ for $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$ fitted via TD Learning. We further observed that the concept of information sharing can be applied to this Q-function learning to reduce variance. This insight led us to develop a new estimator, which we call DQ-IS, taking advantage of two orthogonal ideas of DQ and IS.
In Appendix 19.2, we empirically implement both estimators for the SST setting. A surprising finding is that the DQ estimator shows minimal bias-variance trade-off for SST. Our empirical results indicate that both DQ and DQ-IS exhibit negligible bias, while their variance is close to that of A/B testing and IS respectively. We believe this phenomenon merits further investigation, as it has implications for both DQ theory and the broader literature on bias-variance trade-offs in RL.
9.3 Orthogonal Applications of Information Sharing
It is important to note that our information sharing approach is actually providing a more efficient way to utilize the observed trajectories by leveraging the local treatment structure. This concept of information sharing is not limited to discounted reward MDPs, but is also compatible with various other RL settings.
Average/Total Reward. Under different definitions of the value function, the information sharing idea is still applicable whenever the new treatment is localized. To be more specific, for the average reward MDPs, we can construct a similar plug-in model-based estimator following the corresponding Bellman Equation, ${\bm{V}}=({\bm{I}}-{\bm{P}})^{\sharp}{\bm{r}}$, where the notation $({\bm{I}}-{\bm{P}})^{\sharp}$ denotes the group inverse of the matrix $({\bm{I}}-{\bm{P}})$. Similarly, for the total reward, we have ${\bm{V}}=({\bm{I}}-{\bm{P}})^{-1}{\bm{r}}$ with ${\bm{P}}$ stands for the transition matrix of non-absorbing states. The estimations of ${\bm{P}}$ and ${\bm{r}}$ can still adopt the information sharing idea to reduce the variance.
Function Approximations. Modern practical reinforcement learning often faces the curse of dimensionality, i.e. a prohibitively large or infinite state space. This may render our previous tabular methods of fitting value functions unrealistic. However, one can resort to any estimation methods designed for function approximation while still leveraging the information sharing idea. For example, consider a parameterized class of value functions $V_{\Theta}=\{V_{{\bm{\theta}}}\colon{\bm{\theta}}\in\Theta\}$, where common function approximations include linear (Tsitsiklis and Van Roy 1996, Bhandari et al. 2021) or neural networks (Cai et al. 2019). Given a trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$, typical TD algorithms solve
[…]
which can be viewed as a parameterized version of Eq. (3). When it comes to A/B testing, it’s then evident that one can split trajectories to fit the value function by solving this optimization problem. Additionally, information can still be shared for local treatments, allowing the construction of richer datasets and improving efficiency.
Episodic Setting. Another commonly used framework is the episodic MDP. In this scenario, multiple independent samples of the MDP can be obtained. For instance, the company in Example 1.1 might assume that certain customer groups share the same transition matrix. This allows for the observation of multiple trajectories, enabling the inference of the value function by aggregating data from these various episodes. Given this context, it’s natural to consider extending the concept of Information Sharing as a more sophisticated approach to data pooling in these scenarios. This extension could potentially lead to more efficient and accurate value function estimation by leveraging similarities across different episodes while accounting for their unique characteristics.
10 Conclusion
This paper systematically explores how treatment structure can be utilized to improve experimentation efficiency in MDPs. We first demonstrate that our A/B testing algorithm is efficient for general policies. Subsequently, we show how information can be shared among tests to reduce variance for local treatment policies, thereby overcoming the established lower bound.
Finally, we want to highlight the trade-off between Robustness and Efficiency. The efficiency gain from information sharing relies heavily on the local structure of the new treatment. If the new local treatment on some states also influences transitions and rewards of the other states, our information sharing approach could introduce bias. In contrast, the inference methods that do not utilize the local structure, as introduced in Section 4, are always robust and valid, though potentially at the cost of reduced efficiency.