### v3

# Chen, Simchi-Levi, and Wang

[…]
Chen, Simchi-Levi, and Wang
[…]
Experimenting with Treatment Locality
[…]
Utilizing randomized experiments to evaluate the effect of short-term treatments on the short-term outcomes has been well understood and become the golden standard in industrial practice. However, as service systems become increasingly dynamical and personalized, much focus is shifting toward maximizing long-term outcomes, such as customer lifetime value, through lifetime exposure to interventions. Our goal is to assess the impact of treatment and control policies on long-term outcomes from relatively short-term observations, such as those generated by A/B testing. A key managerial observation is that many practical treatments are local, affecting only targeted states while leaving other parts of the policy unchanged. This paper rigorously investigates whether and how such locality can be exploited to improve estimation of long-term effects in Markov Decision Processes (MDPs), a fundamental model of dynamic systems. We first develop optimal inference techniques for general A/B testing in MDPs and establish corresponding efficiency bounds. We then propose methods to harness the localized structure by sharing information on the non-targeted states. Our new estimator can achieve a linear reduction with the number of test arms for a major part of the variance without sacrificing unbiasedness. It also matches a tighter variance lower bound that accounts for locality. Furthermore, we extend our framework to a broad class of differentiable estimators, which encompasses many widely used approaches in practice. We show that all such estimators can benefit from variance reduction through information sharing without increasing their bias. Together, these results provide both theoretical foundations and practical tools for conducting efficient experiments in dynamic service systems with local treatments.
[…]
Customer Lifetime Value, Randomized Experiments, Local Treatment
1 Introduction
Randomized experiments, commonly known as A/B testing, have emerged as an indispensable tool, enabling organizations across various sectors—including service, retail, manufacturing, and healthcare—to empirically test and refine their strategies for enhancing customer engagement and satisfaction (Imbens and Rubin 2015, Kohavi and Thomke 2017, Kohavi et al. 2020, Koning et al. 2022). In particular, A/B testing is well-established and widely used to compare short-term or immediate outcomes (e.g., whether a customer clicks on an advertisement) resulting from short-term treatments (e.g., displaying a new version of the advertisement).
However, as service systems grow more complex and dynamic, the focus is increasingly shifting from immediate or short-term outcomes to long-term effects on customers or users who are continually exposed to new interventions. For example, online retailers may employ strategies such as regular coupon distribution or cash back, aiming to trade-off short-term profit reductions for potential life-time revenue gains (Simester et al. 2006). Similarly, user interface updates on web platforms tend to have continual exposure to customers after their launches, thus impacting user experience and engagement over extended periods (Huang et al. 2023, Maystre et al. 2023). These practical scenarios require estimating the impact of long-term treatments on long-term outcomes, such as customer lifetime value, using relatively short-term observations from randomized experiments.
In this work, we investigate the randomized experiments on Markov Decision Processes (MDPs), which are widely used to model dynamic service systems over extended time horizons including customer relationships (Pfeifer and Carraway 2000), inventory control (Gong and Simchi-Levi 2023), two-sided platforms (Johari et al. 2022), mobile health (Zhang et al. 2023) and many others. Additionally, as modern systems become increasingly complex with numerous states, the practical challenges of implementing entirely new global policies intensify. Such policies could be overly aggressive, in terms of posing significant risks and requiring substantial infrastructure changes. Consequently, many treatments in practice are becoming more fine-grained and localized. This means that the new treatments may only apply on a small number of targeted states locally. For instance, in the customer relationships, the customer’s states are referring to her recent behaviors or other contextual factors. An online marketplace or service platform might offer a “welcome-back” coupon to the customer only when she has not engaged with the service for an extended period (e.g., Sephora’s Welcome Back Event Dooley 2020). If she is a daily active user, she might not receive it. Similarly, physical retail locations like pharmacies and markets may provide incentives like cash back to the customer only when she makes in-store purchases (e.g., Bed Bath & Beyond Mohammed 2015). Likewise, a typical intervention for recommender systems may involve precise exposure, only taking effect when user click into certain pages such as banners or author profiles, without altering other system components throughout the iterative recommendation process (e.g., Spotify Maystre et al. 2023 and Douyin). These localized treatments are strategic, usually aiming to optimize the lifetime value of customers by tailoring experiences and offers to individual needs and histories. Given the prevalence of local treatments in practice, our research goal is to explore the potential of leveraging treatment locality to optimize data utilization for efficient A/B testing on long-term effects within an MDP environment.
1.1 An Illustrative Example: Customer Lifetime Value
To be more concrete about our research question, we further introduce an application example on improving customer lifetime value, adapted from a seminal work in the marketing literature by Pfeifer and Carraway (2000) and closely related to customer retention and churn management (Hamilton and Singal 2023, Lemmens and Gupta 2020, Kanoria et al. 2024). This example will be revisited throughout the paper.
Example 1.1 (Customer Lifetime Value)
Drawing from the classic Recency, Frequency, Monetary value (RFM) framework, it is often posited by companies that the probability a customer will purchase at the end of any period is solely a function of her recency—the number of periods since her last purchase. Pfeifer and Carraway (2000) model recency as the state space of an MDP and characterize customer lifetime value via the value function, which is the discounted sum of revenue collected in the long-run.
Currently, in order to encourage repeat business and build customer loyalty, the company is exploring the impact of issuing coupons to customers with high recency (i.e., inactive customers at state $s^{5}$) to attract them back. The coupon treatment is considered local, affecting only the state transitions and immediate reward at state $s^{5}$, while leaving all other states unchanged before and after the treatment. Our study can then enable the company to assess how such a coupon policy might affect customer lifetime value in the long run.
1.2 Overview of Main Results
To demonstrate our results, we start with what is arguably the most fundamental and illustrative scenario: the new local treatment available in only one specific state. We refer to such a local treatment as Single-State Treatment (SST). In this setup, the treatment policy of interest applies the new SST whenever possible and adheres to the control policy in all other states. Our objective is to estimate and infer the Average Treatment Effect (ATE), which quantifies the difference in long-term cumulative rewards between the treatment policy and the control policy. We summarize and highlight our main results in detail below.
First, we design an efficient estimator of ATE in general A/B testing on MDPs, where the treatment does not need to be local. Specifically, we show that the two principal approaches from the reinforcement learning (RL) literature, the model-based estimation and temporal difference (TD) learning, remain effective for estimating ATE if the data are split correctly. Our proposed estimator is asymptotically normal and unbiased. Moreover, we show that no unbiased estimator can achieve a smaller variance without considering treatment locality. This result builds on a new efficiency proof framework via the constrained Cramér-Rao bound.
We then introduce an estimator that incorporates the SST structure by sharing data collected from states where the new treatment is not applied. While data sharing naturally increases the sample size for both the treatment and control policies, it also introduces additional correlations that complicate the overall variance. Fortunately, we are able to show that this estimator can always achieve a smaller variance compared to the original estimator that does not share data. Our analysis further reveals that the variance reduction arises specifically from the component of the variance associated with the states that share data. Importantly, when extending to multiple test arms, we illustrate that a major part of the asymptotic variance can be reduced linearly with the number of test arms compared to the estimator that does not use the SST structure. Additionally, we derive another variance lower bound for all unbiased estimators even with the knowledge of the SST, which confirms the optimality of our proposed estimator.
While we have focused on the unbiased estimator for a specific ATE, decision-makers can have many other design choices for A/B testing in dynamic systems. For example, in practice, a slight bias may be acceptable if it substantially reduces variance. Motivated by this, we define a broad class of estimators, called differentiable estimators, which includes many existing estimators in the literature (e.g., Farias et al. 2022, 2023, Huang et al. 2023). We show that information sharing universally reduces variance without increasing bias for all differentiable estimators. Consequently, any proposed differentiable estimators, regardless of its specific design objective, could immediately benefit from the variance reduction through information sharing if the treatment is local. When there are multiple testing arms, a major component of the variance can also get a linear reduction in the number of arms. These results further illustrate the power of treatment locality.
Having demonstrated the variance reduction achieved through SST, we also expand our framework and generalize the results to cover more general local treatments, which are applicable to multiple states, broadening the scope and applicability of our results. We also test the effectiveness of our design on a synthetic environment as well as a Sepsis simulator based on real-world data. Finally, we reveal that under SST, the treatment effect remains consistently positive (or negative) for all states before and after the intervention. This indicates that if the new treatment benefits one state, it benefits all others as well.
1.3 Related Works
Structure-aware Experiment Design. Experimental design originated in economics with basically model-free methods. In recent years, researchers in operations research and management science instead realize that there are many useful structures in operations scenarios to understand, analyze and improve experimentation. Bojinov et al. (2023), Hu and Wager (2022), Ni et al. (2023), Jia et al. (2023), Xiong et al. (2024) consider the switchback design under various scenarios, from temporal interference to spatial interference, from stationary environment to non-stationary environment, and from theory driven perspective to data driven perspective. Glynn et al. (2020), Munro et al. (2024), Wager and Kuang (2021), Bajari et al. (2021), Johari et al. (2022), Zhu et al. (2024), Weng et al. (2024) study the experiments happening on two-sided markets and a common practice is to model the system as a continuous-time Markov chain. Chen et al. (2025a) investigate the A/B testing in e-commerce where the product stock-out can introduce extra bias. Chen et al. (2025b) analyze the bias of experiments for inventory control policies. Li et al. (2015) leverage the sparsity structure of cross-product demand elasticities to greatly reduce the number of experiments required. Li et al. (2023) carefully take the queuing structure into account and propose efficient estimators under the stochastic congestion. Xiong et al. (2023) investigate multi-unit settings with varying treatment starting times. Zhan et al. (2021, 2023) study the post-experiment inference for contextual bandits, which does not capture the state transition as our MDPs. Simester et al. (2020) pioneer at considering treatment structures and suggest comparing outcomes only for customers exposed to different treatment and control actions to improve efficiency empirically. Compared to theirs, we further incorporate complex transition dynamics and provide theoretical guarantees of estimations. The long-term effect with the long-term treatment under the short-term measurement is recently studied by Huang et al. (2023). This paradigm of structure-aware experiment design has witnessed great success in theory and practice in above works with either significant reduction of bias or substantial improvement of efficiency.
Experimenting on MDPs with General Treatments. Experimenting on MDPs is more relevant to our work. Many works adopt MDPs to capture the dynamical system and investigate the long-term outcomes (Pfeifer and Carraway 2000, Murphy 2003, Simester et al. 2006, Tran et al. 2023, Shi et al. 2020, 2023, Farias et al. 2022, 2023), where most previous works consider general treatment, i.e. a treatment will be posed to every state of MDP. The seminal work by Farias et al. (2022) studies the average reward setting. They prove the asymptotic lower bound for unbiased inference on MDPs and show that this lower bound can sometimes grow exponentially with the number of states. They also construct an LSTD(0)-type OPE estimator that is efficient and achieves this lower bound. In our work, we show that similar phenomena can also be observed in the discounted reward settings. Moreover, Farias et al. (2022) study the bias of a Naive estimator and propose an alternative estimator that achieves a notable bias-variance trade-off with an upper bound of asymptotic variance. In the sequel, we demonstrate all three estimators studied in Farias et al. (2022) fall into a general class of differentiable estimator and can thus all benefit from the variance reduction leveraging treatment locality. Finally, some recent work like Shi et al. (2023), Zhang et al. (2023) considers asymptotic results with linear function approximation. In contrast, this paper focuses on a fundamental split of the experimental data and is therefore orthogonal to previous literature on estimator design.
In addition, there is another line of work including Syrgkanis and Zhan (2023), Cheikhi and Russo (2023), Shi et al. (2020), Khamaru et al. (2021), Tran et al. (2023) that considers asymptotic inference on MDPs but under episodic setting or with the assumption of a generative model, which means the MDP can be independently simulated many times. In contrast, we consider estimations that can only sample one trajectory from a non-episodic MDP, which restricts our access to the data and thus imposes additional challenges.
1.4 Structures and Notations
In Section 2, we formally introduce the model and the formulation, introducing the MDP, the treatment policy and the control policy, and the average treatment effect that we care about. Section 2.1 presents the estimation methods for the long-term cumulative rewards of a single, specific policy. In Section 3, we discuss the idea of classical A/B testing in our setting and how to estimate the long-term cumulative rewards for the treatment and control policies from A/B testing. Section 4 provides a new estimator based on the idea of information sharing for the local treatment. Section 5 generalizes the analysis to the general local treatments. Section 6 introduces the differentiable estimators, which can always benefit from the information sharing. In Section 7, we provide some numerical results verifying the idea of information sharing including the real-world sepsis simulation results. Section 8 covers the discussion on some useful properties of the local treatment as well as future work. We conclude our paper in Section 9.
We denote for any $N\in\mathbb{Z}^{+},[N]=\{1,2,\cdots,N\}$. For any vector $x$ and matrix $\bm{A}$}, we define the semi-norm induced by the matrix $A$ as $\|x\|_{\bm{A}}\coloneqq\sqrt{x^{\top}{\bm{A}}x}$. Additionally, for any function on the state space $f\colon{\mathcal{S}}\to\mathbb{R}$, we denote $\operatorname{Diag}_{s\in{\mathcal{S}}}(f(s))\coloneqq\operatorname{Diag}(f(s^{1}),\cdots,f(s^{K}))\in\mathbb{R}^{K\times K}$. Moreover, we define $a\vee b\coloneqq\max\{a,b\}$ and $a\land b\coloneqq\min\{a,b\}$.
2 Model
We start with the general definition of experimenting on MDPs and then specify the Single-State Treatment (SST), which serves as a simple yet illustrative example of local treatment. In Section 5, we generalize this concept to broader local treatment scenarios.
Consider a (non-episodic, finite-state, infinite-horizon) MDP with state space ${\mathcal{S}}:=\{s^{i}\}_{i\in[K]}$ with cardinality $K$. At time $t\in\mathbb{N}$, we denote the state of the MDP by $s_{t}\in{\mathcal{S}}$. Associated with each state is a set of available actions, $\mathcal{A}$}, which influence the state transitions through the transition probability function $P:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\rightarrow[0,1]$. Here, the next state $s_{t+1}$ follows the distribution $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. For each state-action pair $(s,a)$, the instant reward $r$ earned is independently generated by a Gaussian distribution [^1] With standard arguments, our results can be extended to the setting where the second moment of the noise is finite., i.e., $r\sim{\mathcal{N}}(r(s,a),\sigma_{r}^{2}(s,a))$, where we assume $\sigma^{2}_{r}(s,a)<\infty$ for all $(s,a)$ pairs. A policy $\pi:{\mathcal{S}}\rightarrow{\mathcal{A}}$ maps states to (random) actions. The value function is defined as the cumulative discounted reward starting from state $s\in{\mathcal{S}}$ under a policy $\pi$,
$V^{\pi}(s)\coloneqq\mathbb{E}\left[{\sum_{i=0}^{\infty}\gamma^{i}r(s_{i},a_{i})\mid\pi,s_{0}=s}\right],$ (1)
where $0\leq\gamma<1$ is a discount factor that reflects the trade-off between instant and future rewards. The value function has been widely applied to quantify long-term outcomes in various fields including marketing (Pfeifer and Carraway 2000, Simester et al. 2006), ride-hailing platforms (Shi et al. 2023), and video streaming service (Tran et al. 2023, Maystre et al. 2023). We further use ${\bm{V}}^{\pi}\in\mathbb{R}^{K}\coloneqq(V^{\pi}(s^{1}),\cdots,V^{\pi}(s^{K}))^{\top}$ to denote value functions of all states under policy $\pi$. Moreover, associated with each $\pi$, ${\bm{r}}^{\pi}\in\mathbb{R}^{K}$ is the expected reward under each state, defined as ${\bm{r}}^{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}r(s,a)\pi(a\mid s)$ for $s\in{\mathcal{S}}$. ${\bm{P}}^{\pi}\in\mathbb{R}^{K\times K}$ is the transition matrix under $\pi$ with ${\bm{P}}^{\pi}(s,s^{\prime})\coloneqq\sum_{a\in{\mathcal{A}}}P(s^{\prime}\mid s,a)\pi(a\mid s)$ for $s,s^{\prime}\in{\mathcal{S}}$.
We now introduce a special class of treatment policies that we call the Single-State Treatment (SST). Specifically, there exists a pre-determined “crucial” state, which without loss of generality, we assume to be $s^{1}$. The treatment of interest is applied exclusively when the system’s state, $s_{t}$, coincides with $s^{1}$. Formally, we assume that ${\mathcal{A}}=\{{\mathfrak{c}},{\mathfrak{t}}\}$ regardless of states, while the action $\mathfrak{t}$}, representing the new treatment, is only applicable in the crucial state $s^{1}$, while $\mathfrak{c}$} denotes the control or default action. This restriction ensures that the treatment is targeted and its effects are isolated to the crucial state.
Building on the definition of SST, we next define two policies and the corresponding treatment effect we aim to assess.
  • The Control Policy $\pi^{\mathfrak{c}}$: This policy does not implement the new intervention, maintaining $\pi^{\mathfrak{c}}(s)\equiv{\mathfrak{c}}$ across all states $s\in{\mathcal{S}}$. This is typically the default or existing policy used by the organization. Denote the associated transition matrix as ${\bm{P}}^{\mathfrak{c}}$, where the entries are exactly $P(\cdot\mid\cdot,{\mathfrak{c}})$. The vector ${\bm{r}}^{\mathfrak{c}}\in\mathbb{R}^{K}$ is the unknown mean reward for each state under the control whose entries are just $r(\cdot,{\mathfrak{c}})$.
  • The Treatment Policy $\pi^{\mathfrak{t}}$: This policy implements the new intervention $\mathfrak{t}$} exclusively whenever the state is $s^{1}$, while defaulting to $\mathfrak{c}$} in all the other states. Formally, $\pi^{\mathfrak{t}}(s^{1})\equiv{\mathfrak{t}}$ and $\pi^{\mathfrak{t}}(s)\equiv{\mathfrak{c}}$ for all $s\in{\mathcal{S}}\setminus s^{1}$. Accordingly, we denote the associated transition matrix and the mean reward vector as ${\bm{P}}^{\mathfrak{t}}$ and ${\bm{r}}^{\mathfrak{t}}$, respectively.
The treatment effect that we are interested in is the difference of the value function under the treatment policy and the control policy. For simplicity, we specially abbreviate the value functions under $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$, $V^{\pi^{\mathfrak{t}}}(s)$ and $V^{\pi^{\mathfrak{c}}}(s)$, as $V^{{\mathfrak{t}}}(s)$ and $V^{{\mathfrak{c}}}(s)$, respectively. Specifically, for a given initial state $s$, the estimand of our interest is the average treatment effect (ATE),
$\Delta(s)=V^{{\mathfrak{t}}}(s)-V^{{\mathfrak{c}}}(s).$
For all the states, we write the treatment effect in a more compact way as ${\bm{\Delta}}\coloneqq{\bm{V}}^{{\mathfrak{t}}}-{\bm{V}}^{{\mathfrak{c}}}$, where ${\bm{V}}^{{\mathfrak{t}}}$ and ${\bm{V}}^{{\mathfrak{c}}}$ are abbreviations of ${\bm{V}}^{\pi^{\mathfrak{t}}}$ and ${\bm{V}}^{\pi^{\mathfrak{c}}}$, respectively. The central research question of our paper is to estimate the ATE, and especially, how the SST structure can be taken into account and utilized.
2.1 Preliminaries: Inference under a Fixed Policy
In this section, we briefly introduce how to perform inference on the value function under a fixed policy $\pi$ which is not a new task in the literature, as an important cornerstone. There are two streams of methods, both of which are reliant on the well-known Bellman Equation, i.e.,
${\bm{V}}^{\pi}={\bm{r}}^{\pi}+\gamma{\bm{P}}^{\pi}{\bm{V}}^{\pi}.$ (2)
The first stream of methods, known as model-based estimation, is detailed in Algorithm 2. Model-based estimation $\widehat{\bm{V}}^{\textnormal{MB}}$ utilizes the reformulation of the Bellman Equation as ${\bm{V}}^{\pi}=({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-1}{\bm{r}}^{\pi}$, where $\bm{I}$} is the identity matrix. This method estimates ${\bm{V}}^{\pi}$ by plugging in the estimations of the MDP parameters including ${\bm{r}}^{\pi}$ and ${\bm{P}}^{\pi}$. Such an estimation is based on the comprehensive understanding of all model parameters, hence it is termed model-based.
The second methodology, referred to as model-free estimation, is outlined in Algorithm 3. It directly estimates the value function without estimating the model’s parameters. Specifically, for a given tuple $(s_{i},a_{i},s_{i+1},r_{i})$, $r_{i}+\gamma V(s_{i+1})$ is an unbiased estimation of $V(s_{i})$ according to the Bellman Equation. Thus, we obtain the estimator $\widehat{\bm{V}}^{\textnormal{TD}}$ following the well-known least square temporal difference (TD) algorithm, which solves
$\min_{\widehat{\bm{V}}}\sum_{s\in{\mathcal{S}}}\left({\sum_{s_{i}=s}r(s_{i},a_{i})+\gamma\widehat{V}(s_{i+1})-\widehat{V}(s_{i})}\right)^{2}.$ (3)
The two inference algorithms are refereed as Algorithm 2 and 3. Despite of the difference in nature, $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ are actually providing the same point estimation of the value function $\bm{V}$} under very mild assumptions.
Proposition 2.1
Let $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ be the respective output of Algorithm 2 and 3 given the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ sampled under policy $\pi$. Assume all states are visited at least once in $\tau$, then it holds $\widehat{\bm{V}}^{\textnormal{MB}}=\widehat{\bm{V}}^{\textnormal{TD}}$.
With such an equivalence, establishing a property for one estimator inherently confirms it for the other. It is important to emphasize that this equivalence does not imply that the model-based approach and TD learning are identical across all scenarios. We are only talking about finite-space and finite-action MDPs without function approximations in this work. There has been a long debate in the RL literature between model-based learning and TD learning, and we refer to the more relevant surveys and textbooks like Moerland et al. (2023) and Sutton and Barto (2018).
Under mild technical conditions, $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ can be shown to enjoy asymptotic normality.
Proposition 2.2
Assume the Markov chain induced by a fixed policy $\pi$ is uniformly ergodic. Given a trajectory under any initial distribution, it holds as $T\to\infty$, $\sqrt{T}(\widehat{\bm{V}}^{\textnormal{MB}}-{\bm{V}}^{\pi})\xrightarrow{d}{\mathcal{N}}({\bm{0}},\Sigma_{\mathrm{MB}})$.
Here uniform ergodicity is a common technical assumption also used in Farias et al. (2022), Hu and Wager (2022), Jones (2004) and will be specified in Appendix 12.1 along with the detailed form of $\Sigma_{\mathrm{MB}}$. While we expected this result to have been established previously in the RL literature, we were unable to locate an exact reference. For completeness, we provide a full proof in Appendix 12.1. In addition, we highlight another important result, the asymptotic efficiency of these two classical estimators, which may be of independent interest to the RL literature. Specifically, $\Sigma_{\mathrm{MB}}$ is essentially the smallest variance that can be obtained by any unbiased estimator. The proof is given in Appendix 13.1.
Proposition 2.3
Let the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ be sampled from MDP under policy $\pi$ with the stationary distribution ${\bm{\mu}}^{\pi}$. Then for any unbiased estimator of value functions with covariance matrix $\Sigma$, it holds that $\Sigma_{\mathrm{MB}}\preceq T\Sigma$.
3 A/B Testing on MDPs
In the previous section, we have established the inference results with the trajectory generated by a fixed policy $\pi$. To infer the ATE, however, it is necessary to simultaneously estimate ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ from a single trajectory. A commonly employed method for this purpose is A/B testing, where, at any given state $s$, a fair coin flip determines whether to apply $\pi^{\mathfrak{t}}$ or $\pi^{\mathfrak{c}}$. For simplicity, we introduce the experimental policy $\pi^{1/2}$, defined such that $\pi^{1/2}(a\mid s)\coloneqq{1}/{2}\cdot\pi^{\mathfrak{t}}(a\mid s)+{1}/{2}\cdot\pi^{\mathfrak{c}}(a\mid s)$ for any $(s,a)$ pair. Assume that the experimenter executes $\pi^{1/2}$ for $T$ epochs and obtains the trajectory $\tau=\left\{{(s_{i},a_{i},s_{i+1},r_{i})}\right\}_{i\in[T]}$. An example of a realized trajectory might include sequences such as:
${\color[rgb]{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{1,.5,0}(s^{3},{\mathfrak{c}},s^{3},r_{2})}\quad{\color[rgb]{0,0,1}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{1,.5,0}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{0,0,1}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
The blue and orange pairs are the observations generated by the treatment policy $\pi^{\mathfrak{t}}$ and the control policy $\pi^{\mathfrak{c}}$, respectively. Given that we are considering the SST, action $\mathfrak{t}$} is exclusively available at state $s^{1}$, while action $\mathfrak{c}$} is the only option in all other states, regardless of the policy applied. Without considering the SST structure—that is, treating $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ as two completely distinct and general policies—a straightforward approach would involve splitting the trajectory $\tau$ into two separate datasets based on the policy implemented at each time step:
${\color[rgb]{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}}=\{{\color[rgb]{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{t}},i\in[T]\}\,,\qquad{\color[rgb]{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}}=\{{\color[rgb]{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{c}},i\in[T]\}\,$,
where $\pi_{i}$ denotes policy used at time $i$. Leveraging the Markov property of the system, unbiased estimations of ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ can be derived from ${\mathcal{D}}^{\mathfrak{t}}$ and ${\mathcal{D}}^{\mathfrak{c}}$ respectively, using either Algorithm 2 or Algorithm 3. The corresponding algorithm is outlined in Algorithm 1. Although the idea is straightforwards, the estimator can perform in our desired way. More specifically, we can establish the asymptotic normality.
0: Policy $\pi^{\mathfrak{t}},\pi^{\mathfrak{c}}$; Experiment horizon $T$; Discount factor $\gamma$.
1: Execute the experiment policy $\pi^{1/2}$ for $T$ epochs and obtain datasets ${\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}=\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{\mathfrak{t}},i\in[T]\}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}=\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{\mathfrak{c}},i\in[T]\}$.
2: Call Algorithm 2 or Algorithm 3 for $\widehat{{\bm{V}}^{\mathfrak{t}}}$ and $\widehat{{\bm{V}}^{\mathfrak{c}}}$ with ${\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}$ respectively.
3:return  $\widehat{{\bm{V}}^{\mathfrak{t}}}-\widehat{{\bm{V}}^{\mathfrak{c}}}\,$.
Algorithm 1 A/B Testing on Markov Decision Process
Theorem 3.1 (Asymptotic Normality of A/B Testing)
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic, it holds that $\sqrt{T}\left({\widehat{\bm{\Delta}}^{\mathrm{AB}}-{\bm{\Delta}}}\right)\xrightarrow{d}{\mathcal{N}}({\bm{0}},\Sigma_{\mathrm{AB}})$ as $T\to\infty$. The detailed form of $\Sigma_{\mathrm{AB}}$ is provided in the appendix (Theorem 12.4).
Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
$\sqrt{T}\left({\widehat{\Delta}^{\mathrm{AB}}(s^{1})-\Delta(s^{1})}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\mathrm{AB}}^{2}(s^{1}))\,$,
where
$\sigma_{\mathrm{AB}}^{2}(s^{1})=2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}(s)}}^{2}\Big{)}+2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}(s)}}^{2}}\right)\,$,
where $\rho^{\pi}_{s^{1}}(s)$ is defined as $\sum_{i=0}^{\infty}\gamma^{i}\mathbb{P}^{\pi}(s_{i}=s\mid s_{0}=s^{1})$. We use $\sigma^{2}_{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}\mathbb{P}^{\pi}(a\mid s)\sigma^{2}(s,a)$ to denote the variance of the instant rewards of state $s$ under policy $\pi$. The term $\mu_{1/2}(s)$ is the stationary probability of state $s$ under the experiment policy $\pi^{1/2}$, and $\Sigma_{p^{\pi}(s)}\coloneqq\operatorname{Diag}({\bm{p}}_{s})-{\bm{p}}_{s}{\bm{p}}_{s}^{\top}$, where ${\bm{p}}_{s}\in\mathbb{R}^{K}$ is the vector of the transition probability from state $s$ to the other states, $(\mathbb{P}^{\pi}(s^{1}\mid s),\cdots,\mathbb{P}^{\pi}(s^{K}\mid s))^{\top}$.
The term $\rho^{\pi}_{s^{1}}(s)$ is also called the discounted state-visitation measure defined and widely used in the RL literature (Agarwal et al. 2022). The asymptotic variance $\sigma_{\mathrm{AB}}^{2}(s^{1})$ is influenced not only by the state $s^{1}$ itself but also by all the other states. This is consistent with our understanding that our objective is the long-term value and the samples are drawn from an MDP. Although the variance depends on all states, we want to specially highlight that $\sigma^{2}_{\mathrm{AB}}(s^{1})$ can be divided into two groups: one associated exclusively with $\pi^{\mathfrak{t}}$ and the other with $\pi^{\mathfrak{c}}$. Each group has $|{\mathcal{S}}|$ distinct terms, where each term corresponds almost exclusively to one state. Furthermore, each term is inversely proportion to the stationary probability $\mu_{1/2}(s)$. This implies that states which are less frequently visited pose greater challenges in sampling, potentially leading to increased variance. Additionally, each term has two parts, $\sigma_{\pi}^{2}(s)$ and $\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}$, which correspond to the (short-term) variance of the instant rewards and the (long-term) variance related to transitions to the other states.
Finally, we want to remind that Theorem 3.1 is applicable for any general $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$. Our approach segments the trajectory solely based on whether the policy $\pi^{\mathfrak{t}}$ or $\pi^{\mathfrak{c}}$ is used at each step, independent of what exactly $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ are. Consequently, this estimator can be readily extended to estimate ATE for a broader range of treatment policies $\pi^{\mathfrak{t}}$, beyond those confined to SST.
3.1 Variance Lower Bound for General A/B Testing on MDPs
In this subsection, we demonstrate that if we ignore the structure of SST, or if the treatment does not adhere to SST principles at all, then the asymptotic variance $\Sigma_{\textrm{AB}}$ given in Theorem 3.1 is almost optimal for any unbiased estimator of the ATE $\bm{\Delta}$}. The proof is delayed to Appendix 13.2.
Theorem 3.2 (Asymptotic Efficiency of A/B Testing)
If running the mixed policy $\pi^{1/2}$ with the initial state distribution ${\bm{\mu}}_{1/2}$, for any unbiased estimator of ATE $\bm{\Delta}$} with covariance matrix $\Sigma$, it holds that $\Sigma_{\mathrm{AB}}\preceq T\Sigma$, where $\Sigma_{\mathrm{AB}}$ is defined in Theorem 3.1.
This lower bound effectively implies that asymptotic variance of A/B testing cannot be further reduced under general treatment pattern. In recent literature, Farias et al. (2022) derive a similar efficiency result for a similar unbiased A/B testing approach on average-reward MDPs. Their ATE is defined as the difference between average rewards for $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$, i.e. ${\bm{\mu}}^{{\mathfrak{t}}^{\top}}{\bm{r}}^{\mathfrak{t}}-{\bm{\mu}}^{{\mathfrak{c}}^{\top}}{\bm{r}}^{\mathfrak{c}}$. Compared to theirs, our study further incorporates reward variance. From a technical perspective, we employs a distinct proof framework via the constrained Cramér-Rao bound.
4 Variance Reduction: Information Sharing between Test Arms
In this section, we explore how the structure of SST can be leveraged to reduce variance. The underlying principle is that, thanks to the Markov property, the state transitions and reward signals of states $s_{t}\neq s^{1}$ can be shared across different test arms. To illustrate this, let us revisit the trajectory in Section 3. Although $(s^{3},{\mathfrak{c}},s^{3},r_{2})$ is generated by $\pi^{\mathfrak{c}}$, it can also be seen as having been generated by $\pi^{\mathfrak{t}}$ since $\pi^{\mathfrak{t}}(s^{3})=\pi^{\mathfrak{c}}(s^{3})={\mathfrak{c}}$. This observation suggests a potential strategy to use such samples for estimating both ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{t}}$. Consequently, we recolor the trajectory with additional purple on the samples that can be shared between the two policies.
${\color[rgb]{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{.75,0,.25}(s^{3},{\mathfrak{c}},s^{3},r_{2})}\quad{\color[rgb]{.75,0,.25}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{.75,0,.25}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{.75,0,.25}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
Accordingly, we can construct two datasets,
${\color[rgb]{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}}=\{{\color[rgb]{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{t}})\}\cup\{{\color[rgb]{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\},$
${\color[rgb]{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}}=\{{\color[rgb]{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:(s_{i},a_{i})=(s^{1},{\mathfrak{c}})\}\cup\{{\color[rgb]{.75,0,.25}(s_{i},a_{i},s_{i+1},r_{i})}:s_{i}\not=s^{1}\}.$
Applying either Algorithm 2 or Algorithm 3 on the datasets ${\mathcal{D}}^{\mathfrak{t}}_{\textrm{IS}}$ and ${\mathcal{D}}^{\mathfrak{c}}_{\textrm{IS}}$ yields two new estimators $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}$ and $\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$ respectively. Subsequently, the estimator for ATE, denoted as $\widehat{\bm{\Delta}}^{\mathrm{IS}}$, is calculated as $\widehat{\bm{V}}^{\mathfrak{t}}_{\textrm{IS}}-\widehat{\bm{V}}^{\mathfrak{c}}_{\textrm{IS}}$. We will always use “IS” to denote information sharing.
To assess whether information sharing can effectively reduce the variance, we first derive the central limit theorem with the information sharing. The full proof is delayed to Appendix 12.4.
Theorem 4.1 (Asymptotic Normality of Information Sharing)
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic, with SST, it holds that $\sqrt{T}\left({\widehat{\bm{\Delta}}^{\mathrm{IS}}-{\bm{\Delta}}}\right)\xrightarrow{d}{\mathcal{N}}({\bm{0}},\Sigma_{\mathrm{IS}})$ as $T\to\infty$. The detailed form of $\Sigma_{\mathrm{IS}}$ is provided in Appendix 12.4. Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
$\sqrt{T}\left({\widehat{\Delta}^{\mathrm{IS}}(s^{1})-\Delta(s^{1})}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\mathrm{IS}}^{2}(s^{1}))\,$,
where
$\sigma_{\mathrm{IS}}^{2}(s^{1})=2$ $\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)$
$\qquad+\sum_{s\neq s^{1}}\frac{1}{\mu_{1/2}(s)}\left({\big{(}\rho^{\mathfrak{t}}_{s^{1}}(s)-\rho^{\mathfrak{c}}_{s^{1}}(s)\big{)}^{2}\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{\mathfrak{t}}-\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)\,.$
4.1 Analysis of Variance Reduction with Two Test Arms
In this subsection, we dive into the comparison between the asymptotic variance of general A/B testing derived in Theorem 3.1 and that of A/B testing with information sharing in Theorem 4.1. For simplicity, we restrict our focus to the crucial state $s^{1}$. Decomposing the variance based on whether the state is the crucial state or not, we can reorganize the asymptotic variance of A/B testing in Theorem 3.1 as
$\sigma_{\mathrm{AB}}^{2}(s^{1})$ $=\underbrace{2\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)}_{\Sigma^{\mathrm{het}}_{\mathrm{AB}}}$
$\qquad+\underbrace{2\sum_{s\neq s^{1}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}\Big{)}+2\sum_{s\neq s^{1}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)}_{\Sigma^{\mathrm{hom}}_{\mathrm{AB}}}\,.$
The $\Sigma^{\mathrm{het}}_{\mathrm{AB}}$ part is the variance directly related to the crucial state $s^{1}$, and the superscript het reflects that state $s^{1}$ is the only state where $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ may be heterogeneous under SST. The $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ part represents the variance associated with all other states $s\neq s^{1}$ where $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ are homogeneous. We want to remind that under SST, the variances $\sigma^{2}_{\mathfrak{t}}(s)$ and $\sigma^{2}_{\mathfrak{c}}(s)$ are the same for all the other state $s$ except the crucial state $s^{1}$.
Similarly, we can decompose the asymptotic variance of information sharing, $\sigma_{\mathrm{IS}}^{2}(s^{1})$, in the same fashion,
$\sigma_{\mathrm{IS}}^{2}(s^{1})$ $=\underbrace{2\left({\frac{\rho^{\mathfrak{t}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\Big{(}\sigma_{\mathfrak{t}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s^{1})}^{2}\Big{)}+\frac{\rho^{\mathfrak{c}}_{s^{1}}(s^{1})^{2}}{\mu_{1/2}(s^{1})}\left({\sigma_{\mathfrak{c}}^{2}(s^{1})+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s^{1})}^{2}}\right)}\right)}_{\Sigma^{\mathrm{het}}_{\mathrm{IS}}}$
$\qquad+\underbrace{\sum_{s\neq s^{1}}\frac{1}{\mu_{1/2}(s)}\left({\left({\rho^{\mathfrak{t}}_{s^{1}}(s)-\rho^{\mathfrak{c}}_{s^{1}}(s)}\right)^{2}\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{\mathfrak{t}}-\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)}_{\Sigma^{\mathrm{hom}}_{\mathrm{IS}}}\,.$
The above decomposition demonstrates clear structures of asymptotic variance of both estimators. Formally, the following Theorem 4.2 compares $\sigma_{\mathrm{AB}}^{2}(s^{1})$ and $\sigma_{\mathrm{IS}}^{2}(s^{1})$.
Theorem 4.2
Consider SST with treatment policy $\pi^{\mathfrak{t}}$ and control policy $\pi^{\mathfrak{c}}$, it holds
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{ and }\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}.$ (4)
Theorem 4.2 reveals that IS achieves at least an equivalent level of performance to the general A/B testing in terms of asymptotic variance. It aligns with our intuition that the variances associated with $s^{1}$ remain unchanged, as information sharing is not feasible in state $s^{1}$ due to the potential differences in the transition matrix and rewards. Consequently, the observed reduction in variance primarily results from the sharing of information across all other states $s\neq s^{1}$. We would like to highlight that, interestingly, each term of the sum in $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ can be directly related to each term of the sum in $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ through two Cauthy-Schwarz inequalities. More specifically, for any $s\neq s^{1}$,
$\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{\mathfrak{t}}-\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}$ $\leq 2\left\|{\gamma\rho^{\mathfrak{t}}_{s^{1}}(s){\bm{V}}^{t}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}+2\left\|{\gamma\rho^{\mathfrak{c}}_{s^{1}}(s){\bm{V}}^{c}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2},$
$(\rho^{\mathfrak{t}}_{s^{1}}(s)-\rho^{\mathfrak{c}}_{s^{1}}(s))^{2}\sigma_{\mathfrak{c}}^{2}(s)$ $\leq 2\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}\sigma_{\mathfrak{t}}^{2}(s)+2\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}\sigma_{\mathfrak{c}}^{2}(s),$
where the sum of LHS over $s\neq s^{1}$ constitutes the $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ and the sum of RHS forms $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. Such a claim leading to the inequality in Eq. (4). To achieve equality in $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$, the two Cauthy-Schwarz inequalities above must be equal for all $s\neq s^{1}$, which seems to be highly restrictive. In fact, Proposition 4.3 presents the necessary condition for the equality to hold.
Proposition 4.3 (Necessary Condition of Equality)
Equality in Eq. (4) is achieved only if the following two conditions are satisfied simultaneously:
  1. 1.
    the rewards are fixed constants, i.e, $\sigma^{2}_{\mathfrak{c}}(s)=0$ for all $s\neq s^{1}$;
  2. 2.
    Either $r(s,a)=0$ for all $s\in{\mathcal{S}},a\in{\mathcal{A}}$ or $\exists\kappa\neq 0$ such that $\rho^{\mathfrak{t}}_{s^{1}}(s)=\kappa\rho^{\mathfrak{c}}_{s^{1}}(s)$ for all $s\neq s^{1}$.
The proof is delayed to Appendix 12.4.1. This proposition effectively implies that for the equality to hold: either the MDP is trivial with fixed zero reward at every state, or the discounted state visitation measure has to satisfy linear dependence before and after the treatment.
More importantly, we want to highlight that if the new treatment $\mathfrak{t}$} is identical to the control $\mathfrak{c}$} such that for all states $s$, $\rho^{\mathfrak{t}}_{s^{1}}(s)=\rho^{\mathfrak{c}}_{s^{1}}(s)$ and ${\bm{V}}^{\mathfrak{t}}={\bm{V}}^{\mathfrak{c}}$, the term $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ will be zero. In contrast, $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ is a value that can be strictly large than zero. This finding is particularly relevant in contexts like A/A testing. Such an observation is somewhat surprising. With information sharing, the amount of data used to estimate the transition probabilities and the rewards can be thought to be doubled, and thus it seems reasonable to conjecture that IS can at most reduce $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$ to a half, i.e., $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\geq\frac{1}{2}\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$. However, with the example of A/A testing, the conjecture may not always hold true, suggesting that $\Sigma^{\mathrm{hom}}_{\mathrm{IS}}$ could sometimes be significantly smaller than $\Sigma^{\mathrm{hom}}_{\mathrm{AB}}$.
4.2 Variance Lower Bound for Experimentation with SST
In this subsection, we show that IS, though intuitive and simple, is the optimal way to harness treatment locality for consistent A/B testing. Specifically, we present the variance lower bound of any unbiased estimator taking the SST structure into account, which reflects the intrinsic difficulty of A/B testing with SST.
Theorem 4.4 (Asymptotic Efficiency of Information-Sharing)
Assume the treatment is SST with two test arms. If running mixed policy $\pi^{1/2}$ with initial distribution ${\bm{\mu}}_{1/2}$, for any unbiased estimator of ATE $\bm{\Delta}$} with covariance matrix $\Sigma$, it holds $\Sigma_{\mathrm{IS}}\preceq T\Sigma$, where $\Sigma_{\mathrm{IS}}$ is defined in Theorem 4.2.
Theorem 4.4 indicates that our estimator $\widehat{\bm{\Delta}}^{\mathrm{IS}}$ is asymptotically efficient. While Theorem 4.2 highlights that IS does not reduce the variance associated with the crucial state $s^{1}$, Theorem 4.4 affirms that this variance is, in fact, optimal and fundamental.
4.3 (More) Power of IS with Multiple Test Arms
In previous sections, we have shown that IS can do at least as well as A/B testing with two test arms. A practical experiment often contains more than two test arms, such as A/B/n Testing (Kohavi et al. 2020, Russac et al. 2021) or bandit experiments (Simchi-Levi and Wang 2023). In this subsection, we generalize our study of the variance reduction of IS to SST with more than two test arms and show that IS becomes even more powerful in reducing the variance comparing with traditional A/B/n testing. For rigor, we start with the definition of SST with $n$ test arms.
Consider an MDP with state space ${\mathcal{S}}\coloneqq\left\{{s^{i}}\right\}_{i\in[K]}$, where $s^{1}$ is still assumed to be the crucial state of treatment. Unlike the previous model in Section 2 with two test arms, the action set ${\mathcal{A}}=\left\{{{\mathfrak{t}}^{j}}\right\}_{j\in[n]}$ now includes $n$ possible actions. Without loss of generality, we let ${\mathfrak{t}}^{1}$ denote the control action. Recall Example 1.1, a plausible experiment might involve the company testing multiple coupon values to determine which one maximizes customer lifetime value. This implies $n$ SST polices $\Pi\coloneqq\left\{{\pi^{i}}\right\}_{i\in[n]}$ with each $\pi^{i}$ satisfying: $\pi^{i}(s^{1})\equiv{\mathfrak{t}}^{i}$ (i.e., $\pi^{i}({\mathfrak{t}}^{i}\mid s^{1})=1$) and $\pi^{i}(s)\equiv{\mathfrak{t}}^{1}$ (i.e., $\pi^{i}({\mathfrak{t}}^{1}\mid s)=1$) for all $s\in{\mathcal{S}}\setminus s^{1}$. Additionally, each SST policy is associated with its own transition matrix ${\bm{P}}^{{\mathfrak{t}}^{i}}$ and mean reward ${\bm{r}}^{{\mathfrak{t}}^{i}}$.
The primary objective of experimentation in such a setting is to understand the long-term outcome differences among various test arms. Formally, for any pair of SST polices $(\pi^{i},\pi^{j})_{i,j\in[n]}$, we define the ATE as $\Delta_{i,j}(s)=V^{i}(s)-V^{j}(s)$, for given state $s\in{\mathcal{S}}$. For simplicity, throughout this subsection, we restrict our interest to $\Delta_{i,j}(s^{1})$, the ATE under the crucial state $s^{1}$. For A/B/n testing, the experimental policy becomes $\pi^{1/n}(a\mid s)\coloneqq\frac{1}{n}\sum_{i\in[n]}\pi^{i}(a\mid s)$ for any $(s,a)$ pair, meaning that we apply each policy in $\Pi$ with equal probability.
After collecting data, the traditional A/B/n testing would split the trajectory according to different polices and obtain dataset ${\mathcal{D}}^{j}_{\mathrm{ABn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):\pi_{i}=\pi^{j},i\in[T]\}$ for each policy $\pi^{j}$. With information sharing, the observations with $s_{i}\neq s^{1}$ among all $n$ test arms can be shared for all SST polices $\pi$. Then, IS yields the corresponding dataset ${\mathcal{D}}^{j}_{\mathrm{ISn}}\coloneqq\{(s_{i},a_{i},s_{i+1},r_{i}):(s_{i},a_{i})\neq(s^{1},{\mathfrak{t}}^{k}),\forall k\neq j,i\in[T]\}$ for each policy $\pi^{j}$. Applying Algorithm 2 or Algorithm 3 to fit ${\bm{V}}^{j}$ with ${\mathcal{D}}^{j}_{\mathrm{ABn}}$ and ${\mathcal{D}}^{j}_{\mathrm{ISn}}$ respectively can produce two estimators $\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The estimators for the treatment effect become $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ABn}}=\widehat{\bm{V}}^{\textnormal{ABn}}_{i}-\widehat{\bm{V}}^{\textnormal{ABn}}_{j}$ and $\widehat{\bm{\Delta}}_{i,j}^{\textnormal{ISn}}=\widehat{\bm{V}}^{\textnormal{ISn}}_{i}-\widehat{\bm{V}}^{\textnormal{ISn}}_{j}$. The following Corollary 4.5 generalizes Theorems 3.1 and 4.1 revealing the asymptotic normality of these two estimators on the crucial state $s^{1}$.
Corollary 4.5 (Asymptotic Normality of A/B/n Testing with(out) IS)
Assuming the Markov Chain induced by the mixed policy $\pi^{1/n}$ is uniformly ergodic, with SST, it holds as $T\to\infty$, for both $\Gamma\in\{\mathrm{ABn},\mathrm{ISn}\}$,
$\sqrt{T}\left({\widehat{\Delta}_{i,j}^{\Gamma}(s^{1})-\Delta_{i,j}(s^{1})}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\Gamma,(i,j)}^{2}(s^{1}))\,$,
where $\sigma_{\Gamma,(i,j)}^{2}(s^{1})$ can be decomposed into two terms, the variance associated with the crucial state $\Sigma_{\Gamma}^{\mathrm{het}}(i,j)$ and the one with all the other states $\Sigma_{\Gamma}^{\mathrm{hom}}(i,j)$.
We ensure that the both two estimators $\widehat{\Delta}_{i,j}^{\mathrm{ABn}}(s^{1})$ and $\widehat{\Delta}_{i,j}^{\mathrm{ISn}}(s^{1})$ enjoy the asymptotic normality, but with different asymptotic variances. Similar as before, the variance can be decomposed according to states. Theorem 4.6 below characterizes the variance reduction of IS for multiple test arms demonstrating the more power that IS has when facing multiple test arms.
Theorem 4.6
Consider SST with $n$ test arms $\left\{{\pi^{i}}\right\}_{i\in[n]}$, it holds for any pair of test arms $(\pi^{i},\pi^{j})$,
$\Sigma^{\mathrm{het}}_{\mathrm{ISn}}(i,j)=\Sigma^{\mathrm{het}}_{\mathrm{ABn}}(i,j)\,,\qquad\Sigma^{\mathrm{hom}}_{\mathrm{ISn}}(i,j)\leq\frac{2}{n}\Sigma^{\mathrm{hom}}_{\mathrm{ABn}}(i,j),$ (5)
where the equality of the second part can hold only if the reward of states $s\in{\mathcal{S}}\setminus s^{1}$ are fixed, i.e. $\sigma_{{\mathfrak{t}}^{1}}^{2}(s)=0$ for all $s\neq s^{1}$.
This result underscores the effectiveness of IS in leveraging the structure of SST to significantly reduce the variance compared to traditional A/B/n testing. Particularly, we demonstrate that the variance associated with all other states $s\neq s^{1}$, where the control policy is always employed, can be reduced at least linearly with the number of test arms. Thus, the more test arms we have, the greater the reduction in variance achieved by IS. When $n=2$, Theorem 4.6 replicates the findings of Theorem 4.2. When $n\geq 3$, Theorem 4.6 consistently ensures a substantially reduced variance. Remarkably, with $n\geq 4$, IS is expected to cut the variance associated with the non-crucial states by at least half. Similar to Theorem 4.2, the variance associated with crucial state $s^{1}$ is the same as A/B/n testing, as information about this state cannot be shared. Additionally, the equality condition in the inequality of Eq. (5) remains as elusive as before, largely because one necessary condition for equality is the absence of noise in the observed rewards.
5 From SST to Local Treatment
In practice, a local treatment policy may affect more than a single state. Consequently, we generalize the SST concept to accommodate Local Treatment. In this expanded framework, a treatment action is applied to a specific segment of the state space rather than a single state. This approach allows for a more flexible and realistic modeling of interventions in complex MDPs. To illustrate, recall Example 1.1. The company now wants to test a new strategy: posting advertisements to customers when their recency is relatively high—for instance, when it exceeds 2. Such a treatment policy is a Local Treatment, as it is applied only to a specific segment ($s^{3},s^{4},s^{5}$) of the state space. The remaining states ($s^{1},s^{2}$), representing lower recency, are not subject to this treatment.
We then establish the model for Local Treatment. Specifically, there exists a pre-determined “crucial” subset ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$. The treatment policy $\pi^{\mathfrak{t}}$ and control policy $\pi^{\mathfrak{c}}$ will only vary at state $s\in{\mathcal{S}}^{\mathfrak{t}}$. Formally, we define
  • The Control Policy $\pi^{\mathfrak{c}}$: This policy maintains $\pi^{\mathfrak{c}}(s)\equiv{\mathfrak{c}}$ across all states $s\in{\mathcal{S}}$, following the same definition as Section 2.
  • The Treatment Policy $\pi^{\mathfrak{t}}$: This policy now implements the new intervention $\mathfrak{t}$} exclusively whenever at the state $s\in{\mathcal{S}}^{\mathfrak{t}}$, while defaulting to $\mathfrak{c}$} in all the other states in ${\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$. In other words, $\pi^{\mathfrak{t}}(s)\equiv{\mathfrak{t}}$ for all $s\in{\mathcal{S}}^{\mathfrak{t}}$ and $\pi^{\mathfrak{t}}(s)\equiv{\mathfrak{c}}$ for all $s\in{\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$.
Notice that when the size of the crucial subset $|{\mathcal{S}}^{\mathfrak{t}}|=1$, we recover the definition of SST. The average treatment effect $\bm{\Delta}$} follows the same definition as section 2.
Having demonstrated the effectiveness of IS for variance reduction under SST, we can now extend these results to Local Treatment scenarios. It can be expected that we are able to establish similar results to Theorem 3.1, 4.1. Compared to SST, the heterogeneous part of the system becomes a subset ${\mathcal{S}}^{\mathfrak{t}}$ instead of the single just state $s^{1}$. As a result, we can only share information of state $s\in{\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$. The following theorem characterizes this extension from SST.
Theorem 5.1 (Asymptotic Normality with Local Treatment)
Assuming the Markov Chain induced by policy $\pi^{\mathfrak{t}},\pi^{1/2}$ are both uniformly ergodic, with Local Treatment on state ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$, it holds as $T\to\infty$, for $\Gamma\in\{\mathrm{AB},\mathrm{IS}\}$ and any state $s\in{\mathcal{S}}$,
$\sqrt{T}\left({\widehat{\Delta}^{\Gamma}(s)-\Delta(s)}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\Gamma}^{2}(s))\,$,
where $\sigma_{\Gamma}^{2}(s)$ can be decomposed according to the state as
$\sigma_{\Gamma}^{2}(s)$ $=\Sigma_{\Gamma}^{\mathrm{het}}+\Sigma_{\Gamma}^{\mathrm{hom}}\,$,
where $\Sigma_{\Gamma}^{\mathrm{het}}$ and $\Sigma_{\Gamma}^{\mathrm{hom}}$ refers to the variance of states ${\mathcal{S}}^{\mathfrak{t}}$ and ${\mathcal{S}}\setminus{\mathcal{S}}^{\mathfrak{t}}$ respectively for $\Gamma\in\{\mathrm{AB},\mathrm{IS}\}$. The full expression of $\sigma_{\mathrm{AB}}^{2}(s),\sigma_{\mathrm{IS}}^{2}(s)$ can be found in Appendix 14.
Recall in SST, $\Sigma_{\Gamma}^{\mathrm{het}}$ is associated with the single crucial state $s^{1}$, while now we generalize it to the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$. For the asymptotic relative efficiency of three estimators, we then have the following generalization of Theorem 4.2.
Theorem 5.2
Consider Local Treatment Effect on ${\mathcal{S}}^{\mathfrak{t}}\subseteq{\mathcal{S}}$ with two test arms $\pi^{\mathfrak{t}},\pi^{\mathfrak{c}}$, it holds
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\quad\textrm{and}\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\leq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}\,.$
Theorem 5.2 replicates Theorem 4.2. The difference is that now we generalize $\Sigma_{\Gamma}^{\mathrm{het}}$ to variance associated with the crucial subset ${\mathcal{S}}^{\mathfrak{t}}$ rather than single crucial state $s^{1}$ in SST. This demonstrates that the effectiveness of IS diminishes as the number of states that can share information (i.e. the homogeneous part) decreases. In the extreme case of a general treatment where ${\mathcal{S}}^{\mathfrak{t}}={\mathcal{S}}$, meaning all states are treated and no states to share information, IS will reduce to A/B testing. Finally, we can derive a similar generalization to local treatment with multiple test arms as Theorem 4.6. We delay the details to the Appendix 14.1.
6 Information Sharing Always Improves Efficiency for Differentiable Estimators
In the previous sections, we have established the optimal asymptotically unbiased estimators for local treatments in MDPs. In practice, however, decision-makers can have many other design choices. For example, a slight bias may be acceptable if it substantially reduces variance. Moreover, beyond long-term cumulative outcomes, practitioners may also be interested in other long-term treatment effects, such as average or point outcomes. This raises a fundamental question: does information sharing always improve performance regardless of the estimating approach or estimand? To address this, we introduce a broad class of estimators, differentiable estimators, which includes many widely used methods and, importantly, always benefits from information sharing.
6.1 Differentiable Estimators
We now define the differential estimators in the vanilla A/B testing without information sharing. For any piece of data we collect $X_{i}=(s_{i},a_{i},s_{i+1},r_{i})$ and $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$, we denote
${\bm{K}}^{a}_{\mathrm{AB}}(X_{i})\coloneqq 2{\bm{E}}_{s_{i},s_{i+1}}\mathbf{1}[a_{i}=a]\,,\qquad{\bm{R}}^{a}_{\mathrm{AB}}(X_{i})\coloneqq 2{\bm{e}}_{s_{i}}r_{i}\mathbf{1}[a_{i}=a]\,$,
where ${\bm{E}}_{s_{i},s_{i+1}}\in\mathbb{R}^{n\times n}$ is a matrix with all entries equal to zero except for a one at the $(s_{i},s_{i+1})$ entry and ${\bm{e}}_{s_{i}}\in\mathbb{R}^{n}$ is the standard basis vector with a one in the $s_{i}$-th position and zeros elsewhere. Using these definitions, we can have the IPW estimators for $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{P}}^{a}$ and $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{r}}^{a}$,
$\overline{{\bm{K}}^{a}_{\mathrm{AB}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{K}}^{a}_{\mathrm{AB}}(X_{i})\,,\qquad\overline{{\bm{R}}^{a}_{\mathrm{AB}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{R}}^{a}_{\mathrm{AB}}(X_{i})\,.$
Definition 6.1 (Differentiable Estimator)
An estimator $\widehat{{\bm{\Delta}}}^{\textrm{AB}}$ is called a differentiable estimator if
$\widehat{{\bm{\Delta}}}^{\textrm{AB}}=f(\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{AB}}})\,$, (6)
where $f$ is any function that is differentiable w.r.t to $\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{AB}}}$. All the differentiable estimators constitute the class $\mathcal{F}$}.
Note that although our previous ATE $\Delta$ is defined as the difference between two value functions, the estimand ${\bm{\Delta}}^{\mathrm{AB}}$ of differentiable estimator $\widehat{{\bm{\Delta}}}^{\mathrm{AB}}$ can be more general. It can also represent, for example, the difference in average rewards, the difference in finite-horizon rewards, or even the difference in advantage functions. Moreover, the function $f$ need not yield unbiased estimators, and we impose no restrictions beyond differentiability.
Furthermore, we highlight that Definition 6.1 and Eq. (6) subsumes many existing and widely used estimators for MDPs, including:
  • Naive estimator of Farias et al. (2022);
  • Difference-in-Q (DQ) estimator of Farias et al. (2022);
  • Off-policy evaluation estimator in Farias et al. (2022);
  • Longitudinal surrogate model of Huang et al. (2023);
  • Our unbiased A/B testing estimators in Algorithm 1.
This demonstrates that our framework provides a unifying perspective on several strands of the recent literature.
Remark 6.2
The fact that our model-based A/B testing estimator in Algorithm 1 belongs to $\mathcal{F}$} has deeper theoretical implications. In fact, every differentiable model-based estimator falls within $\mathcal{F}$}. More concretely, suppose the MDP is parameterized by $({\bm{P}}^{\mathfrak{t}},{\bm{P}}^{\mathfrak{c}},{\bm{r}}^{\mathfrak{t}},{\bm{r}}^{\mathfrak{c}})$. A differentiable model-based estimator first obtains the estimators $(\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}})$ via Algorithm 2, and then computes the estimate of ATE as $\widehat{{\bm{\Delta}}}^{\textrm{AB}}=g(\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}})$, where $g$ is some functional differentiable w.r.t $\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}}$. We denote the class of all differentiable model-based estimators as $\mathcal{M}$}. The following proposition establishes the inclusion $\mathcal{M}$$\subseteq{\mathcal{F}}$ by explicitly bridging $\widehat{{\bm{P}}}^{\mathfrak{t}},\widehat{{\bm{P}}}^{\mathfrak{c}},\widehat{{\bm{r}}}^{\mathfrak{t}},\widehat{{\bm{r}}}^{\mathfrak{c}}$ to $\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{AB}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{AB}}}$.
Proposition 6.3
Given the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ sampled under policy $\pi$ and assume all states are visited at least once in $\tau$, then it holds $\mathcal{M}$$\subseteq{\mathcal{F}}$.
Proof 6.4
[…]
6.2 Variance Reduction for Differentiable Estimators
Applying the information-sharing idea to differentiable estimators, we again observe a linear reduction in variance for the non-treated states, without introducing additional bias. For clarity, we focus on the SST case, and the extension to general local treatments follows naturally. In particular, information sharing modifies the estimation of model parameters at every state except $s^{1}$. In particular, for $X_{i}=(s_{i},a_{i},s_{i+1},r_{i})$ and $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$, we have
${\bm{K}}^{a}_{\mathrm{IS}}(X_{i})\coloneqq\bigg{(}2\mathbf{1}[a_{i}=a,s_{i}=s^{1}]+\mathbf{1}[s_{i}\neq s^{1}]\bigg{)}{\bm{E}}_{s_{i},s_{i+1}}\,,\quad{\bm{R}}^{a}_{\mathrm{IS}}(X_{i})\coloneqq\bigg{(}2\mathbf{1}[a_{i}=a,s_{i}=s^{1}]+\mathbf{1}[s_{i}\neq s^{1}]\bigg{)}{\bm{e}}_{s_{i}}r_{i}\,.$
And similarly, we denote $\overline{{\bm{K}}^{a}_{\mathrm{IS}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{K}}^{a}(X_{i})$ and $\overline{{\bm{R}}^{a}_{\mathrm{IS}}}\coloneqq\frac{1}{T}\sum_{i=1}^{T}{\bm{R}}^{a}(X_{i})$. Finally, we plug in the estimators into the same $f$ and obtain our information sharing estimator,
$\widehat{{\bm{\Delta}}}^{\textrm{IS}}=f(\overline{{\bm{K}}^{\mathfrak{t}}_{\mathrm{IS}}},\overline{{\bm{K}}^{\mathfrak{c}}_{\mathrm{IS}}},\overline{{\bm{R}}^{\mathfrak{t}}_{\mathrm{IS}}},\overline{{\bm{R}}^{\mathfrak{c}}_{\mathrm{IS}}})\,.$
An immediate observation is that both $\overline{{\bm{K}}^{a}_{\mathrm{IS}}}$ and $\overline{{\bm{K}}^{a}_{\mathrm{AB}}}$ converge almost surely to $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{P}}^{a}$ for $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$. Similarly, $\overline{{\bm{R}}^{a}_{\mathrm{IS}}}$ and $\overline{{\bm{R}}^{a}_{\mathrm{AB}}}$ converge almost surely to $\operatorname{Diag}({\bm{\mu}}_{1/2}){\bm{r}}^{a}$. Since $f$ is differentiable (and thus continuous), ${\bm{K}}^{a}_{\mathrm{AB}}(X_{i})$ and ${\bm{K}}^{a}_{\mathrm{IS}}(X_{i})$ have exactly the same asymptotic bias. More formally, and of greater importance, we have the following variance analysis.
Theorem 6.5
For any differentiable estimators $\widehat{{\bm{\Delta}}}^{\Gamma}=f(\overline{{\bm{K}}^{\mathfrak{t}}_{\Gamma}},\overline{{\bm{K}}^{\mathfrak{c}}_{\Gamma}},\overline{{\bm{R}}^{\mathfrak{t}}_{\Gamma}},\overline{{\bm{R}}^{\mathfrak{c}}_{\Gamma}})$ where $\Gamma\in\{\mathrm{AB},\mathrm{IS}\}$, $\widehat{{\bm{\Delta}}}^{\mathrm{AB}}$ and $\widehat{{\bm{\Delta}}}^{\mathrm{IS}}$ are both asymptotic normal with the same asymptotic bias. Furthermore, their asymptotic variance can be decomposed as $\Sigma_{\Gamma}=\Sigma^{\mathrm{het}}_{\Gamma}+\Sigma^{\mathrm{hom}}_{\Gamma}+\Sigma^{\mathrm{cov}}_{\Gamma}$ where
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}=\Sigma^{\mathrm{het}}_{\mathrm{AB}}\,,\quad\Sigma^{\mathrm{cov}}_{\mathrm{IS}}=\Sigma^{\mathrm{cov}}_{\mathrm{AB}}\,,\textrm{ and }\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}\preceq\Sigma^{\mathrm{hom}}_{\mathrm{AB}}.$
Following Theorem 4.2, the term $\Sigma^{\mathrm{het}}_{\Gamma}$ captures the variance associated with the non-treated states, while $\Sigma^{\mathrm{hom}}_{\Gamma}$ reflects the variance of the state targeted by the SST. The new term introduced here is $\Sigma^{\mathrm{cov}}_{\Gamma}$, which represents the covariance across time. Because the system operates under an MDP, future states depend on the past states and actions. Compared with Theorem 4.2, we can learn that selecting an appropriate functional form of $f$ can eliminate temporal correlation.
For general estimators to A/B/n testing, we can also rigorously show that information sharing can linearly reduce part of the variance. The asymptotic variance can be similarly decomposed as $\Sigma_{\Gamma}(i,j)=\Sigma^{\mathrm{het}}_{\Gamma}(i,j)+\Sigma^{\mathrm{hom}}_{\Gamma}(i,j)+\Sigma^{\mathrm{cov}}_{\Gamma}(i,j)$ where
$\Sigma^{\mathrm{het}}_{\mathrm{IS}}(i,j)=\Sigma^{\mathrm{het}}_{\mathrm{AB}}(i,j)\,,\quad\Sigma^{\mathrm{cov}}_{\mathrm{IS}}(i,j)=\Sigma^{\mathrm{cov}}_{\mathrm{AB}}(i,j)\,,\textrm{ and }\quad\Sigma^{\mathrm{hom}}_{\mathrm{IS}}(i,j)\preceq\frac{2}{n}\Sigma^{\mathrm{hom}}_{\mathrm{AB}}(i,j).$
We defer the formal statement and proofs to Appendix 12.2.
7 Simulations
7.1 Synthetic Data: Customer Lifetime Value
We revisit the example of customer lifetime value (Example 1.1), with a synthetic transition matrix detailed in Appendix 18.1. We assume that under the control policy, when the customer makes a purchase, the company will receive a net income of $30$ while the income in all other states is 0. Specifically, the transition matrix and the reward under the control policy are set as follows.
${\bm{P}}^{\mathfrak{c}}=\begin{bmatrix}0.6&0.4&&&\\ 0.5&&0.5&&\\ 0.4&&&0.6&\\ 0.3&&&&0.7\\ 0.1&&&&0.9\end{bmatrix}\,,\qquad{\bm{r}}^{\mathfrak{c}}=\begin{bmatrix}30\\ 0\\ 0\\ 0\\ 0\end{bmatrix}\,$.
We consider the following two types of treatment:
  1. 1.
    SST. The company considers to offer coupons when the customer reaches the highest recency (recency = 5). We evaluate five different candidate coupon values, each associated with a different treatment effect. These treatments lead to a decrease in reward at state $s^{5}$ by values of $2,2.5,3,3.5,4$, respectively. Correspondingly, they increase the probability of transitioning to state $s^{1}$ by $0.10,0.15,0.20,0.25,0.30$, respectively.
  2. 2.
    Local Treatment. The company plans to offer coupons whenever the customer’s recency is high (recency>>2). We still consider five different candidate values. They uniformly decrease the rewards under states $s^{3},s^{4},s^{5}$ by $2,2.5,3,3.5,4$, respectively and increase the probability transiting to state $s^{1}$ by $0.10,0.11,0.12,0.13,0.14$, respectively.
We additionally introduce another benchmark for comparison, called perfect information (PI), where the decision-makers have the access to the perfect information of the control arm. The PI estimators can always serve as the lower bound of the variance (i.e., the necessary variance). The detailed description of the PI estimators and their properties are provided in Appendix 15.
We run experiments for each candidate value under both cases treatment possible values, aiming to identify the optimal coupon value which are 4 and 2. In each experiment, we run all three estimators over 1,000 independent trajectories and record their empirical mean and standard error. The experiment horizon is adjusted so that the IS estimator’s standard error is smaller than its empirical mean. In Figure 1, we present the simulation results.
Refer to caption Refer to caption
Figure 1: Synthetic simulations for Example 1.1. Diamonds and caps represent the empirical mean and standard error respectively for each estimator. Left: the result for SST policy with $n=5$ test arms and experiment horizon $T=350$. Right: the result for Local Treatment policy with $n=5$ test arms and experiment horizon $T=2000$.
Unbiasedness across all estimators: The empirical means for all three estimators in both experiments demonstrate negligible bias. This aligns with our theoretical results of the consistency of all three estimators.
Variance reduction by IS: In both examples, we observe a significant reduction in variance by at least half through the application of IS. The variance of IS is closer to the variance provided by PI. This illustrates the effectiveness of IS in overcoming the lower bound of original A/B testing by utilizing the local structure. Also, the more states the information can be shared, the closer the variance of IS estimator is to the PI estimator.
While ATEs in both experiments are close, the local treatment case requires a substantially longer horizon for the standard error of IS estimator to decay comparing with the SST case. This discrepancy aligns with the intuition that the SST allows information sharing for more states than the local treatment case. This observation showcases that the efficacy of IS increases with the number of shared states.
7.2 A Real-World Sepsis Simulator
To evaluate our method in a more realistic setting, we employ a simulator of sepsis treatment for intensive care unit (ICU) patients (Komorowski et al. 2018, Oberst and Sontag 2019, Tran et al. 2023). Our focus is on experimenting with various combinations of treatments applicable to patients in critical conditions. The primary objective is to understand the long-term effects of these treatment strategies.
The Environment: The physiological status of patients can be captured by a underlying MDP. The state space encompasses discrete measurements of vital signs (heart rate, blood pressure, oxygen concentration, and glucose levels), categorized into levels (e.g., low, normal, high), along with records of previous treatments. The size of the state space size 720. The reward function assigns values of -1, 0, or +1 based on the number of abnormal vital signs exhibited by a patient. In the absence of treatment intervention, patient signals fluctuate according to specified transition probabilities, reflecting the dynamic nature of the health condition.
The Local Treatment: When patients enter a critical state—characterized by all signals being abnormal—three treatment options become available: antibiotics, vasopressors, and mechanical ventilation. This results in $n=8$ test arms, corresponding to the possible combinations of these treatments. The application of these interventions alters the transition probabilities of the various signals. In our experimental setting, the total number of treated states amounts to 128, representing approximately 18% of the overall state space.
The Simulation: Unlike the previous synthetic example, real-world constraints limit our ability to conduct extended experiments on one individual patient. Instead, we adopt a common practice in sepsis studies: assuming patients with similar contexts follow the same MDP. We pool observation data from a group of non-diabetic patients to estimate the average long-term effect of treatments. Our simulator generates observed trajectories of length 100 for each patient. These patients are initialized according to a state distribution learned a priori.
We run the simulations for $N=250$ patients over $100$ independent runs and then demonstrate the inference results of each estimators. We additionally introduce two more benchmarks, the Naive and Difference-in-Q (DQ) estimators, from Farias et al. (2022), which is originally designed for the general treatment policies on MDPs with two test arms. For the Naive estimators, we naively take the average on the rewards collected under the treatment policy and the control policy, ignoring the interference of state transitions. The DQ estimator strikes a balance between bias and variance under general and global treatment policies in MDPs. In Appendix 18.2, we generalize the DQ theory and provide the expression of DQ estimator for multiple test arms.
Refer to caption Refer to caption
Figure 2: Sepsis Simulator. Diamonds and caps represent the empirical mean and standard error respectively for each estimator. Left: Estimated ATE the best-performing arm among $n=8$ tested arms at patients $N=250$. Right: MSE of each estimator vs. NN the number of patients. IS, PI dominate over all experiment horizon.
Figure 2 summarizes the simulation results as well as the Mean Square Error (MSE) curve with the number of patients. We have the following observations.
Variance Reduction of IS: We observe a significant reduction in variance through the application of IS while still obtaining negligible bias comparing with the A/B testing estimator. The variance of IS is even close to the PI estimator. The IS and PI estimators also enjoy fast decay of MSE compared to all other estimators. This confirms the effectiveness of IS and demonstrates the power of leveraging treatment structure to obtain striking efficiency improvement.
Unbiasedness of IS: Although the Naive estimator has smaller variance than IS, its bias is non-negligible. Furthermore, the Naive estimator struggles to exclude the null effect, potentially resulting in insufficient statistical power. In contrast, the IS and PI estimators are centered around the true ATE.
DQ exhibits bias-variance trade-offs: DQ demonstrates a notable bias-variance trade-off when compared to A/B testing and Naive estimators. DQ achieves variance reduction at the cost of introducing small bias, resulting in a faster decay of MSE. However, DQ fails to leverage the inherent local structures of the treatments, leading to significantly higher MSE compared to IS and PI. As we have discussed that DQ belongs to the general differentiable estimators, we can expect that Is can improve performance over the original DQ formulation.
8 Discussion
8.1 Fairness and Welfare: The Monotonicity of SST
MDPs are complex systems with interconnected states. Therefore, out of fairness and welfare concern, it would be the best for the decision-makers to ensure that a new treatment improves long-term outcomes for customers across all states, rather than benefiting some at the expense of others. Fortunately, we have a theorem that guarantees the “monotonicity” of SST policies.
Theorem 8.1 (Monotonicity of Single-State Treatment Effect)
Consider SST policy $\pi^{\mathfrak{t}}$ with crucial state $s^{1}$. If there exists $s\in{\mathcal{S}}$ such that $\Delta(s)>0$, then for all $s\neq s^{1}$,
$0\leq\Delta(s)<\gamma\Delta(s^{1})\,.$
Vice versa, if there exists $s\in{\mathcal{S}}$ such that $\Delta(s)<0$, then $\gamma\Delta(s^{1})<\Delta(s)\leq 0$ for all $s\neq s^{1}\,.$
This ensures that the treatment effect remains consistently positive (or negative) for all states before and after the intervention. Furthermore, when the treatment effect is negative, all states except the crucial state will experience a welfare loss strictly less than $\gamma\Delta(s^{1})$, where $\gamma\in[0,1)$ is the discount factor. Notably, this monotonicity theorem holds true regardless of the underlying Markov chain’s structure.
Technically, we provide two different proofs for Theorem 8.1 in Appendix 16.2. The first proof is based on an important observation that the treatment effect $\bm{\Delta}$} is an eigenvector of a special matrix associated with the transition probability under the control policy. The second proof offers another perspective to characterize $\bm{\Delta}$} using the advantage function, derived from the well-knwon Performance Difference Lemma (Kakade and Langford 2002, Agarwal et al. 2022) in RL theory. We believe these techniques may be of independent interest to the RL literature.
8.2 Other Applications of Information Sharing
While this paper has mainly focused on A/B testing on a non-episodic MDP under the tabular setting, we believe it’s interesting directions to extend our framework to various other RL settings.
Function Approximations. Modern practical reinforcement learning often faces the curse of dimensionality, i.e. a prohibitively large or infinite state space. This may render our previous tabular methods of fitting value functions unrealistic. However, one can resort to any estimation methods designed for function approximation while still leveraging the information sharing idea. For example, consider a parameterized class of value functions $V_{\Theta}=\{V_{{\bm{\theta}}}\colon{\bm{\theta}}\in\Theta\}$, where common function approximations include linear (Tsitsiklis and Van Roy 1996, Bhandari et al. 2021) or neural networks (Cai et al. 2019). Given a trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$, typical TD algorithms solve
[…]
which can be viewed as a parameterized version of Eq. (3). When it comes to A/B testing, it’s then evident that one can split trajectories to fit the value function by solving this optimization problem. Additionally, information can still be shared for local treatments, allowing the construction of richer datasets and improving efficiency.
Episodic Setting. Another commonly used framework is the episodic MDP. In this scenario, multiple independent samples of the MDP can be obtained. In section 7.2, each patients can be viewed as one such independent MDP. We thus empirically demonstrate that IS is also effective for episodic MDPs and brings about substantial variance reduction. We leave it to future work to derive rigorous theoretical guarantees for information sharing under episodic MDPs.
Policy Optimization. Most RL literature on finding optimal policy (e.g. Azar et al. 2017, Jin et al. 2018, Li et al. 2024) does not impose structural assumption on the action set. Given our modeling of treatment locality, a natural question is whether this structure can also be leveraged to improve the sample efficiency for policy optimization.
9 Concluding Remarks
This paper systematically explores how treatment structure can be utilized to improve the estimation of life-time effects. By exploiting treatment locality, our proposed estimators achieve substantial variance reduction, linear in the number of test arms, without sacrificing unbiasedness. We further show that a broad class of differentiable estimators can universally benefit from information sharing. These results not only establish new theoretical benchmarks but also highlight the practical value of local interventions for efficient experimentation in dynamic systems.
Finally, we want to highlight the trade-off between Robustness and Efficiency. The efficiency gain from information sharing relies heavily on the local structure of the new treatment. If the new local treatment on some states also influences transitions and rewards of the other states, our information sharing approach could introduce bias. In contrast, the inference methods that do not utilize the local structure, as introduced in Section 3, are always robust and valid, though potentially at the cost of reduced efficiency.